### v2

# Estimating invertible processes in Hilbert spaces, with applications to functional ARMA processes

Abstract
Invertible processes are central to functional time series analysis, making the estimation of their defining operators a key problem. While asymptotic error bounds have been established for specific ARMA models on $L^{2}[0,1]$, a general theoretical framework has not yet been considered. This paper fills in this gap by deriving consistent estimators for the operators characterizing the invertible representation of a functional time series with white noise innovations in a general separable Hilbert space. Under mild conditions covering a broad class of functional time series, we establish explicit asymptotic error bounds, with rates determined by operator smoothness and eigenvalue decay. These results further provide consistency-rate estimates for operators in Hilbert space-valued causal linear processes, including functional MA, AR, and ARMA models of arbitrary order.
MSC 2020 subject classifications: 47B38, 60G10, 62F12
Keywords: ARMA; functional time series; invertible processes; linear processes
1 Introduction
Over the past three decades, the field of functional data analysis (fDA) has grown significantly, driven by increasing interest in high-dimensional data obtained from continuous observations across time, space, and frequency; see Horváth and Kokoszka (2012) and Ramsay and Silverman (2005) for textbook length reviews. In many cases, functional data are collected over time, such as continuous electricity price observations transformed into daily curves (Liebl, 2013). Such data are known as functional time series (fTS). Recent advances in their analysis and modeling are discussed in Bosq (2000) and Kokoszka and Reimherr (2017). A typical setup in fTS analysis (fTSA) assumes data reside in a separable, infinite-dimensional Hilbert space $\mathcal{H}$, which we consider throughout. Similar to scalar time series, linear models are most commonly used in fTSA, including $\mathcal{H}$-valued autoregressive (fAR) processes, functional moving average (fMA), and fARMA processes. Under regularity conditions, such processes admit stationary solutions, often represented as $\mathcal{H}$-valued linear and invertible processes. For testing stationarity in general fTS, see Horváth et al. (2014); Aue and van Delft (2020), and van Delft et al. (2021). Invertible representations of stationary processes also arise in non-linear fTS models, such as $\mathcal{H}$-valued volatility processes within f(G)ARCH models, see Hörmann et al. (2013); Aue et al. (2017); Cerovecki et al. (2019); Kühnert (2020). Early works such as Bosq (2000) proposed consistent estimators for fAR operators. Limit theorems for $\mathcal{H}$-valued linear processes are available in Merlevède (1996); Merlevède et al. (1997); Düker (2018), and Račkauskas and Suquet (2010). Aue and Klepsch (2017) and Klepsch and Klüppelberg (2017) focused on point-wise estimation of operators in the linear and inverted representations, and Kühnert (2020) estimated operators of invertible linear processes originating from fGARCH models in $L^{2}[0,1]$. For fAR operators, Caponera and Panaretos (2022) derived explicit rates under mild conditions, while Kara Terki and Mourid (2024) established exponential bounds and convergence rates using sieve estimates.
The fMA and fARMA cases have received less attention. Turbillon et al. (2008) estimated the fMA$(1)$ operator assuming commutativity, and Kuenzer (2024) derived consistent estimators for fARMA operators using a two-step approach, involving sequential AR and MA estimation with dimension reduction via principal components. Central limit theorems for fAR operator estimates are derived in Mas and Pumo (2009).
This article formulates consistency results with quantifiable rates for infinite-dimensional operators defining general $\mathcal{H}$-valued invertible processes with white noise innovations, reyling on Yule–Walker equations and a Tikhonov-regularized estimator, avoiding assumptions tied to specific models. The consistency results are established under a Sobolev-type condition and assuming general consistency rates for certain lagged covariance operator estimates that hold in many standard situations, such as under $L^{p}$-$m$-approximability, a weak dependence concept for fTSA introduced in Hörmann and Kokoszka (2010). We apply these findings to derive consistency for the infinite-dimensional operators defining causal linear processes, including fAR, fMA, and fARMA processes with arbitrary AR and MA orders. Notably, our results are derived under mild conditions, permitting multi-dimensional eigenspaces of the covariance operator, and without requiring specific rates of eigenvalue decay. Provided the operators exhibit a sufficiently high degree of smoothness and the eigenvalue decay of the error covariance operators is slow, our approach yields the near-parametric rate $N^{-1/2}.$
The paper is organized as follows: Section 2 introduces notation and fundamental concepts. Section 3 derives estimators for operators of invertible processes. Sections 45 apply these results to obtain consistent estimators for operators in linear and fARMA processes. Section 6 concludes, with proofs in Appendix A.
2 Preliminaries
2.1 Notation
The additive identity in linear spaces is denoted by 00, and the identity map by $\mathbb{I}$. In Cartesian product spaces $V^{n}$, $n\in\mathbb{N}$, scalar multiplication and addition are defined component-wise. Let $(\mathcal{B},\|\cdot\|)$ be a Banach space and $(\mathcal{H},\langle\cdot,\cdot\rangle)$, $(\mathcal{H}_{\star},\langle\cdot,\cdot\rangle_{\star})$ Hilbert spaces with their respective norms and inner products. Then, $\mathcal{B}^{n}$ with norm $\|x\|^{2}=\sum_{i=1}^{n}\|x_{i}\|^{2}$ is a Banach space, and $\mathcal{H}^{n}$ with inner product $\langle x,y\rangle=\sum_{i=1}^{n}\langle x_{i},y_{i}\rangle$ is a Hilbert space. The spaces of bounded linear, Hilbert–Schmidt (H-S), and nuclear (trace-class) operators $A\colon\mathcal{H}\to\mathcal{H}_{\star}$ are denoted by $\mathcal{L}_{\mathcal{H},\mathcal{H}_{\star}}$, $\mathcal{S}_{\mathcal{H},\mathcal{H}_{\star}}$, and $\mathcal{N}_{\mathcal{H},\mathcal{H}_{\star}}$, respectively, with norms $\|\cdot\|_{\mathcal{L}}$, $\|\cdot\|_{\mathcal{S}}$, and $\|\cdot\|_{\mathcal{N}}$, and H-S inner product $\langle\cdot,\cdot\rangle_{\mathcal{S}}$. For $\mathcal{T}\in\{\mathcal{L,S,N}\}$, set $\mathcal{T_{H}}\coloneqq\mathcal{T}_{\mathcal{H,H}}$. The projection onto the closed subspace spanned by $a_{m},\dots,a_{n}$ of a complete orthonormal system (CONS) $(a_{j})_{j\in\mathbb{N}}$ in $\mathcal{H}$ is denoted $\coprod^{a_{n}}_{a_{m}}\colon\mathcal{H}\to\mathrm{span}\{a_{m},\dots,a_{n}\}\subset\mathcal{H}$. All random variables are defined on a common probability space $(\Omega,\mathfrak{A},\mathbb{P})$. For a sequence $(c_{n})_{n\in\mathbb{N}}\subset(0,\infty)$, $X_{n}=\mathrm{o}_{\mathbb{P}}(c_{n})$ and $X_{n}=\mathrm{O}_{\mathbb{P}}(c_{n})$ as $n\to\infty$ denote $X_{n}/c_{n}$ converges in probability and is stochastically bounded, respectively. The space of random variables $X\in\mathcal{H}$ with $\mathbb{E}\|X\|^{p}<\infty,$ $p\in[1,\infty)$, is denoted by $L^{p}_{\mathcal{H}}=L^{p}_{\mathcal{H}}(\Omega,\mathfrak{A},\mathbb{P})$. A sequence $(X_{k})\subset L^{1}_{\mathcal{H}}$ is centered if $\mathbb{E}(X_{k})=0$ for all $k$ with expectation understood as a Bochner integral. For sequences $(a_{n})_{n\in\mathbb{N}},(b_{n})_{n\in\mathbb{N}}\subset\mathbb{R}$, $a_{n}\asymp b_{n}$ as $n\to\infty$ indicates asymptotic equivalence up to constant factors. Finally, $\mathds{1}_{A}(x)$ denotes the indicator function of a set $A,$ and $\delta_{ij}$ the Kronecker delta. Unless stated otherwise, asymptotic statements refer to $N\to\infty$.
2.2 Weak stationarity and lagged (cross-)covariance operators
Let $(\mathcal{H},\langle\cdot,\cdot\rangle)$ and $(\mathcal{H}_{\star},\langle\cdot,\cdot\rangle_{\star})$ be separable Hilbert spaces with norms $\|\cdot\|=\langle\cdot,\cdot\rangle^{1/2}$ and $\|\cdot\|_{\star}=\langle\cdot,\cdot\rangle_{\star}^{1/2}$, respectively. The cross-covariance operator of two random variables $X\in L^{2}_{\mathcal{H}}$ and $Y\in L^{2}_{\mathcal{H}_{\star}}$ is defined by $\mathscr{C}_{\!X,Y}\coloneqq\mathbb{E}[X-\mathbb{E}(X)]\otimes[Y-\mathbb{E}(Y)],$ where $x\otimes y\coloneqq\langle x,\cdot\rangle y$ denotes the tensor product. The operator $\mathscr{C}_{\!X,Y}\in\mathcal{N}_{\mathcal{H},\mathcal{H}_{\star}}$ has the adjoint $\mathscr{C}^{\ast}_{\!X,Y}=\mathscr{C}_{Y,X}\in\mathcal{N}_{\mathcal{H}_{\star},\mathcal{H}}$, and
$\|\mathscr{C}_{\!X,Y}\|_{\mathcal{N}}\leq\operatorname{\mathds{E}}\!\|X-\operatorname{\mathds{E}}(X)\|\|Y-\operatorname{\mathds{E}}(Y)\|_{\star}\,.$ (2.1)
The covariance operator of $X\in L^{2}_{\mathcal{H}}$ is $\mathscr{C}_{\!X}\coloneqq\mathscr{C}_{\!X,X}.$ It is positive semi-definite, self-adjoint, and
$\|\mathscr{C}_{\!X}\|_{\mathcal{N}}=\operatorname{\mathds{E}}\!\|X-\operatorname{\mathds{E}}(X)\|^{2}.$ (2.2)
[…]
3 Estimation of the operators in invertible processes
This section develops an estimation procedure for operators of invertible processes $(X_{k})_{k\in\mathbb{Z}}$ in a separable Hilbert space $(\mathcal{H},\langle\cdot,\cdot\rangle).$ While such processes are well-defined for bounded linear operators, our approach requires H-S operators. Throughout, we assume the following.
Assumption 3.1.
$\boldsymbol{\varepsilon}=(\varepsilon_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a WN with injective covariance operator $\mathscr{C}_{\boldsymbol{\varepsilon}}\colon\mathcal{H}\rightarrow\mathcal{H}.$
We call a centered process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ invertible if it is weakly stationary and satisfies
$X_{k}=\sum^{\infty}_{j=1}\psi_{j}(X_{k-j})+\varepsilon_{k},\quad k\in\mathbb{Z},$ (3.1)
where $(\varepsilon_{k})_{k\in\mathbb{Z}}$ is a WN, and $\psi_{j}\in\mathcal{L_{H}}$ are bounded linear operators with $\sum^{\infty}_{j=1}\|\psi_{j}\|_{\mathcal{L}}<\infty,$ ensuring the series converges a.s. and in $L^{2}_{\mathcal{H}}.$ Such processes arise in various contexts, including invertible linear processes (Bosq, 2000, Chapter 7), encompassing fAR, as well as fMA and fARMA models under natural conditions (Spangenberg, 2013; Kuenzer, 2024). Moreover, estimating fAR(MA) parameters in invertible processes is relevant for f(G)ARCH models, as their point-wise squared counterparts correspond to fAR(MA) processes (Hörmann et al., 2013; Aue et al., 2017; Kühnert, 2020) with weak WN innovations.
Assumption 3.2.
$(X_{k})_{k\in\mathbb{Z}}\subset L^{2}_{\mathcal{H}}$ is a centered, invertible process with representation (3.1), and the operators $\psi_{j}\in\mathcal{S}_{\mathcal{H}}$.
Consistent estimates for the operators $\psi_{j}$ are derived based on a sample $X_{1},\dots,X_{N}$ from ${\boldsymbol{X}}=(X_{k})$. The proposed estimators are based on lagged (cross-)covariance operators of processes
$X^{[L]}_{k}\coloneqq(X_{k},X_{k-1},\dots,X_{k-L+1})^{\top},\quad k\in\mathbb{Z},$ (3.2)
in the Cartesian product Hilbert space $\mathcal{H}^{L}$. The lag parameter $L=L_{N}\in\mathbb{N}$ is assumed to grow with the sample size. We define the empirical covariance operator $\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}$ of $\boldsymbol{X}^{[L]}=(X^{[L]}_{k})_{k}$ and the empirical lag-1 cross-covariance operator $\hat{\mathscr{D}}_{\!\boldsymbol{X}^{[L]},\boldsymbol{X}}$ of $\boldsymbol{X}^{[L]}$ and $\boldsymbol{X}$ by
$\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}$ $\coloneqq\frac{1}{N}\sum_{k=L}^{N}X^{[L]}_{k}\otimes X^{[L]}_{k},\quad$ $L\leq N,$ (3.3)
$\hat{\mathscr{D}}_{\!\boldsymbol{X}^{[L]},\boldsymbol{X}}$ $\coloneqq\frac{1}{N}\sum_{k=L}^{N-1}X^{[L]}_{k}\otimes X_{k+1},\quad$ $L<N.$ (3.4)
To establish consistency for the estimates for $\psi_{j}$, we assume the following.
Assumption 3.3.
For some sequence $L=L_{N}\rightarrow\infty$ with $L=\mathrm{o}(N^{1/3})$ holds
$\|\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}-\mathscr{C}_{\!\boldsymbol{X}^{[L]}}\|_{\mathcal{S}}$ $=\mathrm{O}_{\operatorname{\mathds{P}}}(L^{3/2}N^{-1/2}),$ (3.5)
$\|\hat{\mathscr{D}}_{\!\boldsymbol{X}^{[L]},\boldsymbol{X}}-\mathscr{D}_{\!\boldsymbol{X}^{[L]},\boldsymbol{X}}\|_{\mathcal{S}}$ $=\mathrm{O}_{\operatorname{\mathds{P}}}(LN^{-1/2}).$ (3.6)
Remark 3.1.
Assumption 3.3 holds for centered processes under $L^{4}$-$m$-approximability (Hörmann and Kokoszka, 2010), as shown in Corollary 3.1 of Kühnert (2024), and for $\pi$-dependence, a notion of weak dependence in metric spaces put forth by Kutta (2025). It also applies to i.i.d. processes with finite fourth moments (Horváth and Kokoszka, 2012). Notably, there are non-causal processes that satisfy Assumption 3.3, such as linear processes in $\mathcal{H}$, that is
$X_{k}=\sum_{\ell=-\infty}^{\infty}\phi_{\ell}(\varepsilon_{k-\ell}),$ (3.7)
where ϕsubscriptitalic-$\phi_{\ell}\in\mathcal{L_{H}}$ with $\sum^{\infty}_{\ell=-\infty}\|\phi_{\ell}\|_{\mathcal{L}}<\infty.$ Then, Assumption 3.3 is satisfied if the linear process operators fulfill the stronger condition $\sum^{\infty}_{k=1}\sum^{\infty}_{\ell=k}\left[\|\phi_{\ell}\|_{\mathcal{L}}+\|\phi_{-\ell}\|_{\mathcal{L}}\right]<\infty.$
The estimates for the operators $\psi_{j}$ in the inverted representation (3.1) are derived based on the approximate Yule–Walker equation
$\mathscr{D}_{\!\boldsymbol{X}^{[L]},\boldsymbol{X}}=\Psi_{\!L}\mathscr{C}_{\!\boldsymbol{X}^{[L]}}+\sum_{\ell>L}\,\psi_{\ell}\mathscr{C}^{1-\ell}_{\!\!\boldsymbol{X}^{[L]},\boldsymbol{X}},$ (3.8)
where the process $\boldsymbol{X}^{[L]}=(X^{[L]}_{k})_{k}\subset\mathcal{H}^{L}$ is defined in (3.2), the operator $\Psi_{\!L}\in\mathcal{L}_{\mathcal{H}^{L}\!,\mathcal{H}}$ is given by
$\Psi_{\!L}\coloneqq(\psi_{1}\,\cdots\,\psi_{L}),$ (3.9)
and the series in (3.8) exists for all fixed $L\in\mathbb{N},$ as it converges in the H-S norm topology under Assumption 3.2, see (A.11). Note that $\Psi_{\!L}$ is identifiable from $\Psi_{\!L}\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$ if $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$ is injective.
Proposition 3.1.
Let Assumptions 3.13.2 hold. Then, $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$ is injective for all $L\in\mathbb{N}.$
Given a sample $X_{1},\dots,X_{N}$ from $\boldsymbol{X}$, we estimate $\Psi_{\!L}$ via Tikhonov regularization, which adds a damping term to the eigenvalues of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$ (see, e.g., Hall and Horowitz, 2005). This yields the estimator
$\hat{\Psi}_{\!L}\coloneqq\hat{\mathscr{D}}_{\!\boldsymbol{X}^{[L]},\boldsymbol{X}}\hat{\mathscr{C}}^{\dagger}_{\!\!\boldsymbol{X}^{[L]}}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}},$ (3.10)
where $\hat{\mathscr{C}}^{\dagger}_{\!\!\boldsymbol{X}^{[L]}}\coloneqq(\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}+\theta_{\!N}\mathbb{I})^{-1}$ is the Tikhonov-type pseudo-inverse of $\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}},$ with identity map $\mathbb{I}$ and positive tuning parameters with $\theta_{\!N}\to 0,$ and $K=K_{N}\to\infty.$ The eigenfunctions $\hat{c}_{1},\dots,\hat{c}_{K}$ correspond to the $K$ largest eigenvalues $\hat{\lambda}_{1}\geq\dots\geq\hat{\lambda}_{K}$ of $\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}$. To analyze the error between $\hat{\Psi}_{\!L}$ and $\Psi_{\!L}$, we decompose $\Psi_{\!L}$ into a finite-dimensional component and a remainder term, controlled via a Sobolev condition similar to that in Hall and Meister (2007). Let $c_{i}=c_{i}(L)$ and $d_{j}$ be the eigenfunctions of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}\colon\mathcal{H}^{L}\to\mathcal{H}^{L}$ and $\mathscr{C}_{\!\boldsymbol{X}}\colon\mathcal{H}\to\mathcal{H}$, respectively. By Assumption 3.2, $\Psi_{\!L}\colon\mathcal{H}^{L}\to\mathcal{H}$ is H-S, and admits the decomposition
$\Psi_{\!L}=\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\Psi_{\!L},c_{i}\otimes d_{j}\rangle_{\mathcal{S}}(c_{i}\otimes d_{j}).$
Assumption 3.4.
For some $\beta>0,$ the limit $S_{\Psi}(\beta)\coloneqq\lim_{L\rightarrow\infty}S_{\Psi_{\!L}}(\beta)<\infty$ exists, where for each $L=L_{N}\in\mathbb{N}$,
$S_{\Psi_{\!L}}\!(\beta)\coloneqq\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\Psi_{\!L},c_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}\,(1+i^{2\beta}\!+j^{2\beta})<\infty.$
Our results bellow allow for repeated eigenvalues $\lambda_{j}=\lambda_{j}(L)$ of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}},$ provided the dimension of each associated eigenspace $\mathcal{V}_{j}(L)$ of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$ is bounded in the following sense.
Assumption 3.5.
There exists a constant $d\in\mathbb{N}$ so that $\sup_{j}\mathrm{dim}(\mathcal{V}_{j}(L))\leq dL.$
In order to state our main result, we define for each $L$ a real-valued sequence $\Lambda_{\ell}=\Lambda_{\ell}(L)$, $\ell\in\mathbb{N}$, related to the reciprocal eigengaps of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$. Let
$\ell_{b}=\inf\left\{j\colon j>\ell,\;\lambda_{j}(L)<\lambda_{\ell}(L)\right\}.$
Define
$\Lambda_{\ell}\coloneqq\frac{1}{\lambda_{\ell}-\lambda_{\ell_{b}}}.$ (3.11)
In other words, $\Lambda_{\ell}$ is the reciprocal eigengap between the eigenspace associated to $c_{\ell}$, and the one after it. Another assumption we need is that the empirical eigenvalues do not occur repeatedly, so that the empirical reciprocal eigengaps are well-defined.
Assumption 3.6.
For all $N$ sufficiently large, with probability one the eigenvalues $\hat{\lambda}_{j}=\hat{\lambda}_{j}(L)$ of $\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}$ satisfy $\hat{\lambda}_{j}\neq\hat{\lambda}_{j+1}$ for all $j\in\{1,...,K\}$ and $L.$
Theorem 3.1.
Let Assumptions 3.13.6 hold. Further, let $K=K_{\!N}\to\infty$ and $\theta_{\!N}\to 0$ be sequences so that for $L=L_{\!N}=\mathrm{o}(N^{1/3})$ from Assumption 3.3 holds $L=\mathrm{O}(K),$ and
$K^{\beta+1/2}\Lambda^{2}_{K}L^{2}=\mathrm{O}(N^{1/2}),\quad\sum_{\ell>L}\|\psi_{\ell}\|_{\mathcal{L}}=\mathrm{O}(\lambda_{K}\Lambda^{2}_{K}K^{1/2}L^{3/2}N^{-1/2}),\quad\theta_{\!N}=\mathrm{O}(\lambda_{K}K^{-\beta}).$
Then,
$\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}=\,\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta}).$
Remark 3.2.
  • (a)
    Theorem 3.1 advances operator estimation for invertible processes by providing explicit rates for the H-S norms of the estimation errors for the entire operators in any separable Hilbert space, allowing multi-dimensional eigenspaces. Previous works, mainly in $\mathcal{H}=L^{2}[0,1]$, often focused on finite-dimensional approximations, omitted rates in infinite dimensions, or assumed one-dimensional eigenspaces. For related results in f(G)ARCH models, see Hörmann et al. (2013); Aue et al. (2017); Cerovecki et al. (2019); Kühnert (2020); and for invertible processes, see Aue and Klepsch (2017); Klepsch and Klüppelberg (2017).
  • (b)
    The rate in Theorem 3.1 is primarily governed by the decay of the eigenvalues $\lambda_{K}=\lambda_{K}(L)$ of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$, the smoothness of $\psi_{j}$ (Assumption 3.4), and, to a lesser extent, the decay of $\|\psi_{j}\|_{\mathcal{L}}$. Estimation is most effective when the $\psi_{j}$ are smooth, and the norms $\|\psi_{j}\|_{\mathcal{L}}$ decay rapidly. Notably, if the eigenvalues of $\psi_{j}$ decay geometrically, the parametric rate $N^{-1/2}$ can be almost reached (see Examples 3.1 and 5.1).
  • (c)
    Assumption 3.4 ensures square-summability of $\|\psi_{j}\|_{\mathcal{S}}$ via Parseval’s identity. Assumption 3.5 imposes minimal restrictions and even accomodates the case that $\boldsymbol{X}=(X_{k})$ is i.i.d. if the eigenspaces of the $\mathscr{C}_{\!\boldsymbol{X}}^{\prime}$ have dimension up to $d,$ since $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}=\mathrm{diag}(\mathscr{C}_{\!\boldsymbol{X}},\dots,\mathscr{C}_{\!\boldsymbol{X}}).$ For conditions on the eigengaps $f(j)=\lambda_{j}-\lambda_{j-1}$, see, e.g., Hall and Horowitz (2007), equation 3.2.
We conclude this section by illustrating the interplay between the conditions in Theorem 3.1.
Example 3.1.
Consider the invertible process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ given by
$X_{k}=\sum_{j=1}^{\infty}\psi^{j}(X_{k-j})+\varepsilon_{k},\quad k\in\mathbb{Z},$
where $(\varepsilon_{k})_{k\in\mathbb{Z}}\subset L^{4}_{\mathcal{H}}$ is an i.i.d. process with injective covariance operator $\mathscr{C}_{\boldsymbol{\varepsilon}}$, having eigenvalues $a_{\ell}$ and eigenfunctions $e_{\ell}$. Suppose the operator $\psi$ admits the spectral representation
$\psi=\sum^{\infty}_{\ell=1}p_{\ell}(e_{\ell}\otimes e_{\ell}),$ (3.12)
with $1/2>p_{1}\geq p_{2}\geq\cdots>0,$ and $p_{\ell}\to 0$. Since $\|\psi\|_{\mathcal{L}}=p_{1}<1/2$, iterative application of the invertible representation yields the linear representation
[…]
Thus, $\mathscr{C}_{\!\boldsymbol{X}}=\mathscr{C}^{0}_{\!\boldsymbol{X}}$ and $\mathscr{C}^{h}_{\!\boldsymbol{X}}$ share eigenfunctions $e_{\ell}$ with $\mathscr{C}_{\!\boldsymbol{\varepsilon}}$, with eigenvalues
$\mu^{0}_{\ell}=\frac{a_{\ell}}{2-4p_{\ell}}(2-3p_{\ell}),\quad\mu^{h}_{\ell}=\frac{a_{\ell}}{2-4p_{\ell}}(1-p_{\ell})(2p_{\ell})^{|h|},\quad h\neq 0.$
The operator-valued matrix $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}=(\mathscr{C}^{j-i}_{\!\boldsymbol{X}})^{L}_{i,j}$ has eigenfunctions
$c_{\ell}=e_{\lfloor(\ell-1)/L\rfloor+1}v_{\ell,((\ell-1)\,\mathrm{mod}\,L)+1},$
with eigenvectors $v_{\ell,1},\dots,v_{\ell,L}$ corresponding to eigenvalues $t_{\ell,1}\geq\dots\geq t_{\ell,L}$ of $T_{\ell}=(\mu^{|j-i|}_{\ell})^{L}_{i,j}.$ By Gershgorin’s theorem,
$\big{|}t_{\ell,m}-\mu^{0}_{\ell}\big{|}\leq\max_{1\leq i\leq L}\sum^{L}_{j=1,j\neq i}\mu^{|j-i|}_{\ell}=\sum^{L-1}_{j=1}\mu^{|j-\lfloor(L+1)/2\rfloor|}_{\ell},\quad m=1,\dots,L.$
Since this sum converges as $L\to\infty$ and is of order $\mathrm{O}(a_{\ell})$ as $\ell\to\infty$ (as is $\mu^{0}_{\ell}$), the eigenvalues of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$ satisfy $\lambda_{\ell}(L)\asymp a_{\lfloor\ell/L\rfloor},$ ensuring Assumption 3.5. For Assumption 3.4, setting $d_{j}=e_{j}$ and $\Psi_{\!L}=(\psi^{1},\dots,\psi^{L})^{\top}$ yields $\langle\Psi_{\!L},c_{i}\otimes d_{j}\rangle_{\mathcal{S}}\leq\delta_{ij}\sum_{m\leq L}p^{m}_{i},$ where $\delta_{ij}$ is the Kronecker delta. Hence, for $\beta<b-1/2$ if $p_{\ell}\asymp\ell^{-b}$ for some $b>1$, or any $\beta>0$ if $p_{\ell}\asymp q^{\ell}$ for some $q\in(0,1),$ it holds that
$\lim_{L\to\infty}S_{\Psi_{\!L}}(\beta)=\lim_{L\to\infty}\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\Psi_{\!L},c_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}(1+i^{2\beta}+j^{2\beta})\leq\sum_{i=1}^{\infty}\,\frac{p^{2}_{i}(1+2i^{2\beta})}{(1-p_{i})^{2}}<\infty.$
It remains to determine the sequences in Theorem 3.1. With $\theta_{\!N}$ chosen freely within the given constraints, we know $\sum_{\ell>L}\|\psi_{\ell}\|_{\mathcal{L}}\propto\|\psi\|^{L}_{\mathcal{L}}=\mathrm{o}(N^{-1/2}),$ whenever $L$ is a positive power of $N$. Valid choices of $K=K_{\!N}$ with $K^{\beta+1/2}\Lambda^{2}_{K}L^{2}\asymp N^{1/2}$ is for given $\beta$ and $L$ determined by the decay of the eigenvalues $\lambda_{j}.$ E.g., for $a_{\ell}\asymp\ell^{-a}$ for some $a>1,$ it holds $\lambda_{K}\asymp(KL^{-1})^{-a}$ and thus $\Lambda_{K}\asymp(KL^{-1})^{a+1}.$ Hence, with $L=L_{\!N}\asymp K^{c}$ for small $c\in(0,1)$ and $K\asymp N^{1/(5+4a(1-c)+2\beta)},$ Theorem 3.1 yields
$\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}\Big{(}N^{-\frac{\beta}{5+4a(1-c)+2\beta}}\Big{)}.$
For e.g., $a_{\ell}\asymp\ell^{-2}$ (the eigenvalue decay rate for Brownian motion errors, see Jaimez and Bonnet, 1987), the rates are, for instance, close to $N^{-1/8}$ for $p_{\ell}\asymp\ell^{-13/6}$, $N^{-1/4}$ for $p_{\ell}\asymp\ell^{-13/2}$, and approximately the parametric rate $N^{-1/2}$ for $p_{\ell}\asymp\rho^{\ell}$, $\rho\in(0,1)$.
4 Consequences for causal linear processes
We now turn to the estimation of operators in the linear representation of invertible, causal linear processes. Let $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ follow a causal linear process (see (3.7) for general cases),
$X_{k}=\sum^{\infty}_{j=0}\phi_{j}(\varepsilon_{k-j}),\quad k\in\mathbb{Z},$ (4.1)
where $(\varepsilon_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a WN, and $\phi_{j}\in\mathcal{L}_{\mathcal{H}}$ are operators with $\phi_{0}\coloneqq\mathbb{I}$ and $\sum_{j=1}^{\infty}\|\phi_{j}\|_{\mathcal{L}}<\infty.$ The process is invertible if it admits a representation as in (3.1). For a comprehensive treatment of linear processes in function spaces, see Bosq (2000), particularly Section 7 for existence and invertibility results. We now focus on estimating the operators $\phi_{j},$ which is crucial for prediction and confidence set construction in $\mathcal{H}$-valued models. The relationship between estimates from an inverted representation of $(X_{k})$ and its causal form follows from the next proposition.
Proposition 4.1.
Let Assumption 3.1 hold. Then the following is true.
  • (a)
    If for some $i\in\mathbb{N}$ it holds that $\sum^{\infty}_{\ell=0}A_{\ell}(\varepsilon_{i-\ell})=\sum^{\infty}_{\ell=0}B_{\ell}(\varepsilon_{i-\ell}),$ then $A_{\ell}=B_{\ell}$ for all $\ell\in\mathbb{N}.$
  • (b)
    If $(X_{k})$ is a causal linear process and if for some $i\in\mathbb{N}$, $\sum^{\infty}_{\ell=0}A_{\ell}(X_{i-\ell})=\sum^{\infty}_{\ell=0}B_{\ell}(X_{i-\ell}),$ then $A_{\ell}=B_{\ell}$ for all $\ell\in\{0,1,2,\dots\}.$
For an invertible linear process $(X_{k})$ in Assumption 3.2 such that
$X_{k}=\sum^{\infty}_{i=1}\psi_{i}(X_{k-i})+\varepsilon_{k},\quad\mbox{ and }\quad X_{k-i}=\sum^{\infty}_{j=1}\phi_{j}(\varepsilon_{k-i-j})+\varepsilon_{k-i},$
for all $i,k\in\mathbb{Z}$, we have due to $\phi_{0}=\mathbb{I}$ that
$X_{k}$ $=\sum^{\infty}_{i=1}\,\Big{[}\sum^{i}_{j=1}\psi_{j}\phi_{i-j}\Big{]}(\varepsilon_{k-i})+\varepsilon_{k},\quad k\in\mathbb{Z}.$
This representation, the linear representation (4.1), and Proposition 4.1 (a) yield
$\phi_{i}=\sum^{i}_{j=1}\psi_{j}\phi_{i-j},\quad i\in\mathbb{N}.$
Hence, with $\hat{\psi}_{j}=\hat{\Psi}^{(j)}_{\!L}$ denoting the $j$’th component of the estimator $\hat{\Psi}_{\!L}$ in (3.10) for $\Psi_{\!L}=(\psi_{1},\dots,\psi_{L}),$ a reasonable estimate for $\phi_{i}$ is $\hat{\phi}_{i},$ defined iteratively with $\hat{\phi}_{0}\coloneqq\mathbb{I},$ and
$\hat{\phi}_{i}\coloneqq\sum^{i}_{j=1}\,\hat{\psi}_{j}\hat{\phi}_{i-j},\quad i\in\mathbb{N}.$ (4.2)
The consistency result for the complete operators in the inverted representation carries over to those in the linear representation.
Proposition 4.2.
Let the conditions of Theorem 3.1 hold, and let $\phi_{i}\in\mathcal{S_{H}}$ for all $i.$ Then,
$\|\hat{\phi}_{i}-\phi_{i}\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}\big{(}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}\big{)}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta}),\quad i\in\mathbb{N}.$
5 Consequences for fARMA processes
The asymptotic results in Section 3 extend to the estimation of fAR, fMA, and fARMA parameters. These models are extensively studied in Bosq (2000), particularly the fAR case, while Spangenberg (2013) provides general conditions for causal and invertible representations in Banach spaces. Estimation results for the fAR(1) operator are available in Bosq (2000) and Caponera and Panaretos (2022). Additionally, Turbillon et al. (2008) estimated the fMA(1) operator assuming commutativity with the error covariance operator, and Aue and Klepsch (2017); Klepsch et al. (2017) studied prediction for fARMA$(p,q)$ processes. Kuenzer (2024) derived consistent fMA(1) and fARMA(1,1) operator estimates using an adaptation of the Hannan–Rissanen method (Hannan and Rissanen, 1982). By using operator estimates from both the linear and inverted representations, we derive consistent fAR, fMA, and fARMA operator estimates with explicit rates.
5.1 fAR processes
A centered process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a functional autoregressive process of order $p\in\mathbb{N}$ (fAR$(p)$) if
$X_{k}=\sum^{p}_{i=1}\,\alpha_{i}(X_{k-i})+\varepsilon_{k},\quad k\in\mathbb{Z},$
where $(\varepsilon_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a WN, and $\alpha_{i}\in\mathcal{L_{H}}$ are operators with $\alpha_{p}\neq 0.$ With $\Psi_{\!p}$ in (3.9), it holds $\Psi_{\!p}=(\alpha_{1}\,\cdots\,\alpha_{p}),$ and the exact Yule–Walker equation $\mathscr{C}^{1}_{\!\boldsymbol{X}^{[p]},\boldsymbol{X}}=\Psi_{\!p}\mathscr{C}_{\!\boldsymbol{X}^{[p]}}$ is fulfilled. Then, with
$\hat{\alpha}_{i}\coloneqq\hat{\Psi}^{(i)}_{\!p},\quad i=1,\dots,p,$
which are the $i$’th components of $\hat{\Psi}_{\!p}$ in (3.10), we obtain the following corollary of Theorem 3.1.
Corollary 5.1.
Let the conditions of Theorem 3.1 with $L=p$ hold, and let $\alpha_{i}\in\mathcal{S_{H}}$ for all $i.$ Then,
$\max_{1\leq i\leq p}\|\hat{\alpha}_{i}-\alpha_{i}\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta}).$
Example 5.1.
Let $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ be the fAR(1) process
$X_{k}=\alpha_{1}(X_{k-1})+\varepsilon_{k},\quad k\in\mathbb{Z},$
where $\boldsymbol{\varepsilon}=(\varepsilon_{k})_{k\in\mathbb{Z}}\subset L^{4}_{\mathcal{H}}$ is i.i.d. with injective covariance operator $\mathscr{C}_{\boldsymbol{\varepsilon}}$ having eigenvalues $a_{\ell}$. The operator $\alpha_{1}$ has the same form as $\psi$ in Example 3.1. Since $\|\alpha_{1}\|_{\mathcal{L}}<1$, Theorems 3.1–3.2 in Bosq (2000) ensure a unique weakly stationary solution, where the covariance operator has the form
$\mathscr{C}_{\!\boldsymbol{X}}=\mathscr{C}_{\boldsymbol{\varepsilon}}\sum_{i=0}^{\infty}\alpha_{1}^{2i},$
with eigenvalues $\lambda_{\ell}\propto a_{\ell}.$ To derive the asymptotic rate, we follow Example 5.1, setting $L=p=1.$ Assumption 3.4 is fulfilled if the eigenvalues $p_{i}$ of $\alpha_{1}$ fulfill $\sum_{i=1}^{\infty}i^{2\beta}p_{i}^{2}<\infty.$ Then, if $a_{\ell}\asymp\ell^{-a}$ for some $a>1$, e.g., Corollary 5.1 ensures that $K$ can be chosen so that
$\|\hat{\alpha}_{1}-\alpha_{1}\|_{\mathcal{S}}=\mathrm{O}_{\mathbb{P}}\Big{(}N^{-\frac{\beta}{5+4a+2\beta}}\Big{)}.$
5.2 fMA processes
A centered process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a functional moving-average process of order $q\in\mathbb{N}$ (fMA$(q)$) if
$X_{k}=\sum^{q}_{j=1}\,\beta_{j}(\varepsilon_{k-j})+\varepsilon_{k},\quad k\in\mathbb{Z},$
where $(\varepsilon_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a WN, and $\beta_{j}\in\mathcal{L_{H}}$ are operators with $\beta_{q}\neq 0.$ Morever, let
$\hat{\beta}_{j}=\hat{\phi}_{j},\quad j=1,\dots,q,$
with $\hat{\phi}_{j}$ in (4.2). Based on the previous considerations, we immediately obtain the following result.
Corollary 5.2.
Let the conditions of Theorem 3.1 hold, and let $\beta_{j}\in\mathcal{S_{H}}$ for all $j.$ Then, it holds
$\max_{1\leq j\leq q}\|\hat{\beta}_{j}-\beta_{j}\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta}).$
5.3 fARMA processes
We finally turn to the estimation of fARMA processes. A centered process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is called a functional autoregressive moving-average process of order $p,q\in\mathbb{N}$ (fARMA$(p,q)$) if
$X_{k}=\sum^{p}_{i=1}\,\alpha_{i}(X_{k-i})+\sum^{q}_{j=1}\,\beta_{j}(\varepsilon_{k-j})+\varepsilon_{k},\quad k\in\mathbb{Z},$ (5.1)
where $(\varepsilon_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a WN, and $\alpha_{i},\beta_{j}\in\mathcal{L_{H}}$ are operators with $\alpha_{p}\neq 0,\beta_{q}\neq 0.$ We combine this fARMA$(p,q)$ representation with its inverted form (3.1). As with multivariate ARMA processes, the operators $\alpha_{i}$ and $\beta_{j}$ are generally not identifiable. Sufficient conditions for identifiability are provided in Kuenzer (2024) (equation (10) and Propositions 1–3). We assume that the fARMA models considered are identifiable and admit invertible, centered, linear process solutions.
To describe the main idea, we first discuss estimation for fARMA(1,1)11(1,1)( 1 , 1 ) processes of the form
$X_{k}=\alpha_{1}(X_{k-1})+\,\beta_{1}(\varepsilon_{k-1})+\varepsilon_{k},\quad k\in\mathbb{Z}.$ (5.2)
From this identity, Proposition 4.1 and $\varepsilon_{k}=X_{k}-\sum^{\infty}_{i=1}\psi_{i}(X_{k-i})$ it follows with $\psi_{0}=\mathbb{I},$
$\psi_{i}=\begin{cases}\,\alpha_{1}+\beta_{1},&\textnormal{if}\;\,i=1,\\ \,-\beta_{1}\psi_{i-1},&\textnormal{if}\;\,i>1.\\ \end{cases}$ (5.3)
Although $\hat{\Psi}_{\!L}$ in (3.10) provides consistent asymptotic estimates for all $\psi_{i}$, consistent estimates for the fARMA(1,1)11(1,1)( 1 , 1 ) operators cannot be directly obtained, as $\psi_{1}\in\mathcal{S_{H}}$ does not possess a bounded inverse. However, using $\hat{\psi}_{i}=\hat{\Psi}^{(i)}_{\!L}$, equation (5.3) leads to the following estimators for $\alpha_{1}$ and $\beta_{1}$:
$\hat{\alpha}_{1}\coloneqq\hat{\psi}_{1}-\hat{\beta}_{1},\qquad\hat{\beta}_{1}\coloneqq-\hat{\psi}_{2}\hat{\psi}_{1}^{\dagger}\coprod^{\hat{f}_{M}}_{\hat{f}_{1}},$ (5.4)
where $\hat{\psi}_{1}^{\dagger}\coloneqq(\hat{\psi}_{1}\hat{\psi}_{1}^{\ast}+\gamma_{N}\mathbb{I})^{-1},$ and $(M_{N})\subset\mathbb{N}$ and $(\gamma_{N})\subset(0,\infty)$ are sequences with $M_{N}\to\infty$ and $\gamma_{N}\to 0$. Additionally, $\hat{f}_{1},\dots,\hat{f}_{M}$ are the eigenfunctions corresponding to the eigenvalues $\hat{\rho}_{1}\geq\dots\geq\hat{\rho}_{M}\geq 0$ of $\hat{\psi}_{1}\hat{\psi}_{1}^{\ast}\colon\mathcal{H}\to\mathcal{H}$. The sequence $(\rho_{j},f_{j})$ represents the eigenpairs of the operator $\psi_{1}\psi_{1}^{\ast}\colon\mathcal{H}\to\mathcal{H}$. The estimates in (5.4) are consistent under the following assumptions, where the first assumption ensures identifiability of $\beta_{1}$ from $\psi_{2}=-\beta_{1}\psi_{1}.$
Assumption 5.1.
The image of the operator $\psi_{1}\colon\mathcal{H}\rightarrow\mathcal{H}$ lies dense.
Assumption 5.2.
The operators $\alpha_{1},\beta_{1}\in\mathcal{S_{H}}$ are H-S, and for some $\gamma>0,$ it holds
$\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\beta_{1},f_{i}\otimes f_{j}\rangle^{2}_{\mathcal{S}}(1+i^{2\gamma}\!+j^{2\gamma})<\infty.$
Assumption 5.3.
The dimensions of all eigenspaces of $\psi_{1}\psi^{\ast}_{1}$ are bounded above by some $d\in\mathbb{N}.$
Assumption 5.4.
For all $N$ sufficiently large, with probability one the eigenvalues of $\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}$ satisfy $\hat{\rho}_{j}\neq\hat{\rho}_{j+1}$ for all $j\in\{1,...,M\}.$
Theorem 5.1.
Suppose $(X_{k})$ is the fARMA(1,1)11(1,1)( 1 , 1 ) process in (5.2). Further, let the conditions of Theorem 3.1 and Assumptions 5.15.4 hold, and let $M=M_{\!N}\rightarrow\infty$ with $M^{\gamma+1/2}P^{2}_{\!M}=\mathrm{O}(K^{\beta})$ and $\gamma_{N}=\mathrm{O}(\rho_{M}M^{-\gamma}),$ with $P_{M}\coloneqq(\rho_{M}-\rho_{M_{b}})^{-1}$ as in (3.11). Then,
$\max\!\big{\{}\|\hat{\alpha}_{1}-\alpha_{1}\|_{\mathcal{S}},\|\hat{\beta}_{1}-\beta_{1}\|_{\mathcal{S}}\big{\}}=\mathrm{O}_{\operatorname{\mathds{P}}}(M^{-\gamma}).$
The procedure for arbitrary $p,q$ is similar as for $p=q=1.$ Similar to the transformations that led to (5.3), we obtain for any $p$ and $q,$ with $\alpha_{i}=\beta_{j}=0$ for $i>p,j>q,$
$\sum^{\infty}_{i=1}\,\psi_{i}(X_{k-i})\,=\sum^{\max(p,q)}_{i=1}\Big{[}\alpha_{i}+\beta_{i}-\sum^{i-1}_{j=1}\,\beta_{j}\psi_{i-j}\Big{]}(X_{k-i})\,-\sum^{\infty}_{i=\max(p,q)+1}\Big{[}\sum^{q}_{j=1}\,\beta_{j}\psi_{i-j}\Big{]}(X_{k-i}).$
Comparing the left- and right-hand sides of the above and applying Proposition 4.1 (b), we see that
$\psi_{i}=\begin{cases}\,\alpha_{i}+\beta_{i}-\sum^{i-1}_{j=1}\,\beta_{j}\psi_{i-j},&\textnormal{if}\;\,1\leq i\leq\max(p,q),\\ \,-\sum^{q}_{j=1}\beta_{j}\psi_{i-j},&\textnormal{if}\;\,i>\max(p,q).\\ \end{cases}$ (5.5)
We proceed to obtain estimates for $\beta_{i}$ for $i>\max(p,q),$ and then for $\alpha_{i}$. Since (5.5) involves sums of compositions of the $\beta_{j}$ with $\psi_{i-j}$ for $i>\max(p,q),$ we cannot immediately retrieve estimates for $\beta_{j}$ by applying Tikhonov regularized versions of the estimates $\psi_{i-j}$. For $i=p+q,$ with
$B_{q}\coloneqq(\beta_{1}\,\cdots\,\beta_{q})\in\,\mathcal{L}_{\mathcal{H}^{q}\!,\mathcal{H}},$
$\Psi^{\prime}_{[i]}\coloneqq(\psi_{p+q+i-1}\,\cdots\,\psi_{p+i})\in\,\mathcal{L}_{\mathcal{H}^{q}\!,\mathcal{H}},\quad i\geq 0,$ (5.6)
the identity (5.5) becomes $\psi_{p+q}=-B_{q}\Psi^{\prime\,\top}_{[0]}.$ Identifiability of $B_{q}$ and thus of $\beta_{1},\dots,\beta_{q}$ is given if the image of $\Psi^{\prime\,\top}_{[0]}\in\,\mathcal{L}_{\mathcal{H},\mathcal{H}^{q}}$ would lie dense. This condition is never met. Nonetheless, identifiability of $B_{q}$ can be established through the following relationship:
$\Psi^{\prime}_{[q]}=-B_{q}\,\boldsymbol{\prod}\;,$ (5.7)
where $\boldsymbol{\prod}\in\mathcal{L}_{\mathcal{H}^{q}}$ is the operator-valued matrix defined by
$\boldsymbol{\prod}\,\coloneqq\begin{bmatrix}\,\Psi^{\prime}_{\![q-1]}\,\\ \,\Psi^{\prime}_{\![q-2]}\,\\ \,\vdots\,\\ \,\Psi^{\prime}_{\![0]}\,\\ \end{bmatrix}=\begin{bmatrix}\,\psi_{p+2q-2}&\psi_{p+2q-3}&\cdots&\psi_{p+q-1}\,\\ \,\psi_{p+2q-3}&\psi_{p+2q-4}&\cdots&\psi_{p+q-2}\,\\ \,\vdots&\vdots&\cdots&\vdots\,\\ \,\psi_{p+q-1}&\psi_{p+q-2}&\cdots&\psi_{p}\\ \end{bmatrix}.$ (5.8)
For $\Psi^{\prime}_{[i]}$ and $\boldsymbol{\prod},$ we use the estimates $\hat{\Psi}^{\prime}_{[i]}$ and $\boldsymbol{\hat{\prod}},$ respectively, which are defined by
$\hat{\Psi}^{\prime}_{[i]}\coloneqq(\hat{\psi}_{p+q+i-1}\,\cdots\,\hat{\psi}_{p+i})\in\,\mathcal{L}_{\mathcal{H}^{q}\!,\mathcal{H}},\quad i\geq 0,\mbox{ and }$ (5.9)
^$\boldsymbol{\hat{\prod}}\,\coloneqq\begin{bmatrix}\,\hat{\Psi}^{\prime}_{\![q-1]}\,\\ \,\hat{\Psi}^{\prime}_{\![q-2]}\,\\ \,\vdots\,\\ \,\hat{\Psi}^{\prime}_{\![0]}\,\\ \end{bmatrix}=\begin{bmatrix}\,\hat{\psi}_{p+2q-2}&\hat{\psi}_{p+2q-3}&\cdots&\hat{\psi}_{p+q-1}\,\\ \,\hat{\psi}_{p+2q-3}&\hat{\psi}_{p+2q-4}&\cdots&\hat{\psi}_{p+q-2}\,\\ \,\vdots&\vdots&\cdots&\vdots\,\\ \,\hat{\psi}_{p+q-1}&\hat{\psi}_{p+q-2}&\cdots&\hat{\psi}_{p}\\ \end{bmatrix}.$ (5.10)
Due to Lemma A.2 (b), the operator-valued vectors $\Psi^{\prime}_{[i]},\hat{\Psi}^{\prime}_{[i]}$ and matrices $\prod,\hat{\prod}$ are under Assumption 3.4 elements of $\mathcal{S}_{\mathcal{H}^{q}\!,\mathcal{H}}$ and $\mathcal{S}_{\mathcal{H}^{q}},$ respectively, thus ,^^$\prod\prod^{\ast},\hat{\prod}\,\hat{\prod}^{\ast}\in\mathcal{N}_{\mathcal{H}^{q}}$ are nuclear. As a result of (5.5) and (5.7), reasonable estimates for $\alpha_{i}$ and $\beta_{j},$ respectively, are
$\hat{\alpha}_{i}\coloneqq\hat{\psi}_{i}+\hat{B}_{[i]}\hat{\Psi}^{\prime\prime\prime\,\top}_{[i]},\;1\leq i\leq p,\qquad\hat{\beta}_{j}\coloneqq\begin{cases}\,\hat{B}^{(j)}_{q},&\textnormal{if}\;\,1\leq j\leq q,\\ \,0,&\textnormal{if}\;\,j>q.\\ \end{cases}$ (5.11)
In these definitions, $\hat{B}_{q}$ stands for the estimator for $B_{q}$ defined by
$\hat{B}_{q}\coloneqq-\hat{\Psi}^{\prime}_{[q]}\,\boldsymbol{\hat{\prod}}^{\dagger}\coprod^{\hat{h}_{\!M}}_{\hat{h}_{1}},$
where ^(^^$\boldsymbol{\hat{\prod}}^{\dagger}\coloneqq(\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}+\gamma_{N}\hskip 0.50003pt\mathbb{I})^{-1},$ $(M_{N})\subset\mathbb{N}$ and $(\gamma_{N})\subset(0,\infty)$ are sequences with $M_{N}\to\infty$ and $\gamma_{N}\rightarrow 0,$ and $\hat{h}_{1},\dots,\hat{h}_{\!M}$ are the eigenfunctions associated to the eigenvalues $\hat{\zeta}_{1}\geq\cdots\geq\hat{\zeta}_{M}\geq 0$ of $\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}.$ Further, $\hat{B}_{[i]},$ $B_{[i]},$ $\hat{\Psi}^{\prime\prime\prime}_{[i]},$ $\Psi^{\prime\prime\prime}_{[i]}\in\mathcal{S}_{\mathcal{H}^{\max(i,q)},\mathcal{H}}$ are for each $i,q$ defined by
$\hat{B}_{[i]}\coloneqq\begin{cases}\,\hat{B}_{i},&\textnormal{if}\;\,1\leq i<q,\\ \,\hat{B}_{q},&\textnormal{if}\;\,i\geq q,\\ \end{cases}\qquad\text{and}\qquad B_{[i]}\coloneqq\begin{cases}\,B_{i},&\textnormal{if}\;\,1\leq i<q,\\ \,B_{q},&\textnormal{if}\;\,i\geq q,\\ \end{cases}$ (5.12)
$\hat{\Psi}^{\prime\prime\prime}_{[i]}\coloneqq\begin{cases}\,\hat{\Psi}^{\prime\prime}_{[0]},&\textnormal{if}\;\,1\leq i\leq q,\\ \,\hat{\Psi}^{\prime\prime}_{[i-q]},&\textnormal{if}\;\,i>q,\\ \end{cases}\qquad\text{and}\qquad\Psi^{\prime\prime\prime}_{[q]}\coloneqq\begin{cases}\,\Psi^{\prime\prime}_{[0]},&\textnormal{if}\;\,1\leq i\leq q,\\ \,\Psi^{\prime\prime}_{[i-q]},&\textnormal{if}\;\,i>q,\\ \end{cases}$ (5.13)
with $\hat{\Psi}^{\prime\prime}_{[i]}$ and $\hat{\Psi}^{\prime}_{[i]}$ matching $\hat{\Psi}^{\prime}_{[i]}$ in (5.9) and $\Psi^{\prime}_{[i]}$ in (5.6) for all $i$, except for a reversed sign in the first component. To derive consistent estimates we make the following assumptions.
Assumption 5.5.
The image of the operator-valued matrix $\boldsymbol{\prod}\colon\mathcal{H}^{q}\rightarrow\mathcal{H}^{q}$ lies dense.
Assumption 5.6.
The operators $\alpha_{1},\dots,\alpha_{p},\beta_{1},\dots,\beta_{q}\in\mathcal{S_{H}}$ are H-S, and for some $\gamma>0,$ it holds
$\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle B_{q},h_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}\,(1+i^{2\gamma}\!+j^{2\gamma})<\infty,$
where $h_{i}$ and $d_{j}$ denote the eigenfunctions of $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}\colon\mathcal{H}^{q}\rightarrow\mathcal{H}^{q}$ and $\mathscr{C}_{\!\boldsymbol{X}}\colon\mathcal{H}\rightarrow\mathcal{H},$ respectively.
Assumption 5.7.
The dimensions of all eigenspaces of $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}$ are bounded above by some $d\in\mathbb{N}.$
Assumption 5.8.
For all $N$ sufficiently large, with probability one the eigenvalues of $\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}$ satisfy $\hat{\zeta}_{j}\neq\hat{\zeta}_{j+1}$ for all $j\in\{1,...,M\}.$
Theorem 5.2.
Suppose $(X_{k})$ is the fARMA$(p,q)$ process in (5.1). Further, let the conditions of Theorem 3.1 and Assumptions 5.55.8 hold, and let $M=M_{\!N}\to\infty$ with $M^{\gamma+1/2}Z^{2}_{\!M}=\mathrm{O}(K^{\beta})$ and $\gamma_{N}=\mathrm{O}(\zeta_{M}M^{-\gamma})$, where $Z_{\!M}\coloneqq(\zeta_{M}-\zeta_{M_{b}})^{-1}$ as in (3.11), with eigenvalues $\zeta_{j}$ of $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}$. Then,
$\max_{1\leq i\leq p,1\leq j\leq q}\!\big{\{}\|\hat{\alpha}_{i}-\alpha_{i}\|_{\mathcal{S}},\|\hat{\beta}_{j}-\beta_{j}\|_{\mathcal{S}}\big{\}}$ $=\mathrm{O}_{\operatorname{\mathds{P}}}(M^{-\gamma}).$
6 Concluding remarks
This article establishes consistent Yule–Walker-type estimators for the operators of invertible processes in separable Hilbert spaces under Sobolev conditions using Tikhonov regularization. We derive consistent estimators for functional linear, AR, MA, and ARMA processes of arbitrary order, leveraging asymptotic consistency results for growing finite-dimensional projections. The convergence rates primarily depend on operator smoothness in the inverted representation and error covariance regularity, with lesser sensitivity to operator norm residuals. Notably, high smoothness give near-parametric rate $N^{-1/2}$ even under rough error covariance operators.
Granger and Andersen (1978) introduced a broader notion of invertibility for real-valued processes with i.i.d. errors, defining $(X_{k})$ as invertible if $\operatorname{\mathds{E}}(\hat{\varepsilon}_{k}-\varepsilon_{k})^{2}\to 0$ as $k\to\infty$, where
$\varepsilon_{k}$ $=X_{k}-f(X_{k-1},\dots,X_{k-p},\varepsilon_{k-1},\dots,\varepsilon_{k-q}),$
$\hat{\varepsilon}_{k}$ $=X_{k}-f(X_{k-1},\dots,X_{k-p},\hat{\varepsilon}_{k-1},\dots,\hat{\varepsilon}_{k-q}),$
with given initial values and i.i.d. centered innovations $\varepsilon_{k}$ with finite second moments. When $f$ is unknown, it is replaced by an estimate. Hallin (1980, 1981) extended this concept by replacing time limits with backward estimates, allowing for time dependence. This notion generalizes to separable Banach spaces $(\mathcal{B},\|\cdot\|)$. A process $(X_{k})\subset\mathcal{B}$ is mildly Granger–Andersen invertible if
$\|\hat{\varepsilon}_{k_{N}}\!-\varepsilon_{k_{N}}\|=\mathrm{o}_{\operatorname{\mathds{P}}}(1),$
where $(\varepsilon_{k})$ is i.i.d., $N$ is the sample size, and $k_{N}\to\infty$. This condition is weaker than standard invertibility (3.1), requiring only convergence in probability without assuming finite second moments or constraining $(k_{N})$. Standard invertibility implies mild Granger–Andersen invertibility, but not vice versa. As noted by Granger and Andersen (1978), certain functional time series, such as bilinear processes in separable Hilbert spaces (Hörmann and Kokoszka, 2010), satisfy mild but not standard invertibility. Others, like the non-linear MA$(1)$ process $X_{k}=\alpha_{1}(\varepsilon^{2}_{k-1})+\varepsilon_{k}$ in $L^{2}[0,1]$, may be stationary yet fail to be invertible in both senses.
A natural direction for future research is extending estimation methods to mild Granger-Andersen invertibility under suitable assumptions. Further, developing a general estimation framework for Banach spaces, potentially even metric spaces, remains an open and promising challenge.
Acknowledgements
The authors are grateful to the two referees and the associate editor for their constructive comments on an earlier version of this article, which greatly improved the quality of the paper. Parts of the paper were written while Sebastian Kühnert was employed at the University of California, Davis.
Funding
Sebastian Kühnert was partially supported by TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (Project number 520388526) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). Gregory Rice was partially supported by the Discovery Grant RGPIN 50503-11525 3100 105 from the Natural Science and Engineering Research Council of Canada.
Appendix A Proofs
Various proofs utilize the operator-valued Hölder inequality for Schatten class operators, along with several other standard results in operator theory (Weidmann, 1980, section 7.1). Let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be Hilbert spaces. Then, a compact linear operator $A\colon\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}$ is a Schatten class operator of order $1\leq p<\infty$ if its singular values $s_{1}(A)\geq s_{2}(A)\geq\dots\geq 0$ are $p$-fold summable. The set of such operators, denoted by $\mathcal{S}_{p,\mathcal{H}_{1},\mathcal{H}_{2}}$ can be equipped with the norm $\|A\|_{p}\coloneqq(\sum^{\infty}_{i=1}s^{p}_{i}(A))^{1/p},$ $A\in\mathcal{S}_{p,\mathcal{H}_{1},\mathcal{H}_{2}}.$ The spaces of nuclear $\mathcal{N}_{\mathcal{H}_{1},\mathcal{H}_{2}}$ and H-S operators $\mathcal{S}_{\mathcal{H}_{1},\mathcal{H}_{2}}$ match $\mathcal{S}_{1,\mathcal{H}_{1},\mathcal{H}_{2}}$ and $\mathcal{S}_{2,\mathcal{H}_{1},\mathcal{H}_{2}},$ respectively. For the space of bounded linear operators $\mathcal{L}_{\mathcal{H}_{1},\mathcal{H}_{2}},$ we write $\mathcal{S}_{\infty,\mathcal{H}_{1},\mathcal{H}_{2}},$ and put $\|\cdot\|_{\infty}\coloneqq\|\cdot\|_{\mathcal{L}}.$ Note that for any $1\leq p<q\leq\infty$ it holds that $\mathcal{S}_{q,\mathcal{H}_{1},\mathcal{H}_{2}}\subsetneq\mathcal{S}_{p,\mathcal{H}_{1},\mathcal{H}_{2}},$ with $\|\cdot\|_{p}\leq\|\cdot\|_{q}.$
Lemma A.1.
Let $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}$ be Hilbert spaces, and let $p,q,r\in[1,\infty]$ with $\frac{1}{p}+\frac{1}{q}=\frac{1}{r},$ and $\frac{1}{\infty}\coloneqq 0.$ Then, for any $A\in\mathcal{S}_{p,\mathcal{H}_{2},\mathcal{H}_{3}}$ and $B\in\mathcal{S}_{q,\mathcal{H}_{1},\mathcal{H}_{2}},$ it holds that $AB\in\mathcal{S}_{r,\mathcal{H}_{1},\mathcal{H}_{3}},$ with
$\|AB\|_{r}\leq\|A\|_{p}\|B\|_{q}.$
Lemma A.2.
Let $\mathcal{H}$ be a Hilbert space, and let $B_{ij}\in\mathcal{L_{H}}$ and $S_{ij}\in\mathcal{S_{H}}$ be bounded and H-S operators, respectively, where $i=1,\dots,m\in\mathbb{N},j=1,\dots,n\in\mathbb{N}.$ Then, the following holds.
  • (a)
    For the operator-valued matrix $\boldsymbol{B}\coloneqq(B_{ij})_{1\leq i\leq m,1\leq j\leq n}$ it holds that $\boldsymbol{B}\in\mathcal{L}_{\mathcal{H}^{m}\!,\mathcal{H}^{n}},$ with
    $\|\boldsymbol{B}\|_{\mathcal{L}}\leq\sum^{m}_{i=1}\|(B_{i1}\;\cdots\;B_{in})\|_{\mathcal{L}}.$
  • (b)
    For the operator-valued matrix $\boldsymbol{S}\coloneqq(S_{ij})_{1\leq i\leq m,1\leq j\leq n}$ it holds that $\boldsymbol{S}\in\mathcal{S}_{\mathcal{H}^{m}\!,\mathcal{H}^{n}},$ with
    $\|\boldsymbol{S}\|^{2}_{\mathcal{S}}=\sum^{m}_{i=1}\sum^{n}_{j=1}\|S_{ij}\|^{2}_{\mathcal{S}}.$
Proof.
See Weidmann (1980), Theorem 7.8. ∎
In several places, we make use of the following upper bounds. From (2.2), the definition of the elements of $\boldsymbol{X}^{[L]}=(X^{[L]}_{k})_{k\in\mathbb{Z}}$ in (3.2), and weak stationarity of $\boldsymbol{X}=(X_{k})_{k\in\mathbb{Z}}$, it follows that
$\|\mathscr{C}_{\!\boldsymbol{X}^{[L]}}\|_{\mathcal{N}}\!=\operatorname{\mathds{E}}\!\|X^{[L]}_{0}\|^{2}=L\operatorname{\mathds{E}}\!\|X_{0}\|^{2}.$ (A.1)
Thus, due to (2.1) and the Cauchy–Schwarz inequality, $\|\mathscr{C}^{h}_{\!\boldsymbol{X}^{[L]}}\|_{\mathcal{N}}\leq\|\mathscr{C}_{\!\boldsymbol{X}^{[L]}}\|_{\mathcal{N}}$ for all $h\in\mathbb{Z},$ and
$\|\mathscr{C}^{h}_{\!\boldsymbol{X}^{[L]},\boldsymbol{X}}\|_{\mathcal{N}}\leq\sqrt{L}\operatorname{\mathds{E}}\!\|X_{0}\|^{2},\quad h\in\mathbb{Z}.$ (A.2)
A.1 Proofs of results in Section 3
Proof of Proposition 3.1.
First, suppose $L=1$. The covariance operator $\mathscr{C}_{\!\boldsymbol{X}^{[1]}}=\mathscr{C}_{\!\boldsymbol{X}}$ of the invertible process in (3.1) can be expressed in terms of the lagged cross-covariance operators $\mathscr{C}^{h}_{\!\boldsymbol{X}}=\mathbb{E}(X_{0}\!\otimes\!X_{h})$ for $h\in\mathbb{Z}$, and the error covariance operator $\mathscr{C}_{\boldsymbol{\varepsilon}}$ as:
$\mathscr{C}_{\!\boldsymbol{X}}=\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\psi_{j}\mathscr{C}^{i-j}_{\!\boldsymbol{X}}\psi^{*}_{i}+\mathscr{C}_{\boldsymbol{\varepsilon}}.$ (A.3)
The series converges in the nuclear norm topology, as the triangle inequality, operator-valued Hölder inequality, $\|\psi^{\ast}_{i}\|_{\mathcal{L}}=\|\psi_{i}\|_{\mathcal{L}},$ $\|\mathscr{C}^{h}_{\!\boldsymbol{X}}\|_{\mathcal{N}}\leq\|\mathscr{C}_{\!\boldsymbol{X}}\|_{\mathcal{N}}$ and Assumption 3.2 imply
$\Big{\|}\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\psi_{j}\mathscr{C}^{i-j}_{\!\boldsymbol{X}}\psi^{\ast}_{i}\Big{\|}_{\mathcal{N}}$ $\leq\|\mathscr{C}_{\!\boldsymbol{X}}\|_{\mathcal{N}}\Big{(}\sum_{i=1}^{\infty}\|\psi_{i}\|_{\mathcal{L}}\Big{)}^{2}<\infty.$
Since $\mathscr{C}_{\boldsymbol{\varepsilon}}$ is positive definite by Assumption 3.1, $\mathscr{C}_{\!\boldsymbol{X}}$ is injective if the double sum in (A.3) is positive semi-definite. In fact, for every $x\in\mathcal{H}$,
$\Big{\langle}\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\psi_{j}\mathscr{C}^{i-j}_{\!\boldsymbol{X}}\psi^{*}_{i}(x),x\Big{\rangle}$ $=\mathbb{E}\Big{(}\sum_{i=1}^{\infty}\,\langle X_{i},\psi^{*}_{i}(x)\rangle\Big{)}^{2}\geq 0.$
Next, let $L\in\mathbb{N}.$ Suppose by way of contradiction that there exists a $v=(v_{1},\dots,v_{L})^{\top}\in\mathcal{H}^{L}$ such that $v\neq 0,$ and
$0=\langle\mathscr{C}_{\!\boldsymbol{X}^{[L]}}(v),v\rangle=\!\mbox{ var}\Big{(}\sum_{i=0}^{{L-1}}\langle X_{-i},v_{i}\rangle\Big{)}.$
If this is so there exists an $r\leq L-1$ so that $v_{r}\neq 0$, and almost surely
$\langle X_{-r},v_{r}\rangle=-\sum_{i=r+1}^{L-1}\langle X_{-i},v_{i}\rangle.$ (A.4)
Notice the right-hand side of the above is measurable with respect to $\sigma(\varepsilon_{i},\;i\leq-r-1)$. Due to the invertible process representation (3.1), the left-hand side of the above is equal to
$\sum^{\infty}_{j=1}\,\langle\psi_{j}(X_{-r-j}),v_{r}\rangle+\langle\varepsilon_{-r},v_{r}\rangle.$
Therefore, multiplying each side of (A.4) with $\langle\varepsilon_{-r},v_{r}\rangle$ and taking expectations gives $\langle\mathscr{C}_{\varepsilon}(v_{r}),v_{r}\rangle=0$, with $v_{r}\neq 0$, which contradicts the injectivity of $\mathscr{C}_{\boldsymbol{\varepsilon}}.$
To prove Theorem 3.1, we first consider the idealized case where $\Psi_{\!L}$ is finite-dimensional (hence H-S) and diagonalizable with respect to the CONS $(c_{i}\otimes d_{j})=(c_{i}(L)\otimes d_{j})_{i,j}$ of $\mathcal{S}_{\mathcal{H}^{L},\mathcal{H}}$, where $(\lambda_{j},c_{j})=(\lambda_{j}(L),c_{j}(L))_{j}$ and $(\mu_{j},d_{j})$ are the eigenpair sequences of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$ and $\mathscr{C}_{\!\boldsymbol{X}}$, respectively.
Assumption A.1.
For $K\in\mathbb{N}$ and all $L\in\mathbb{N}$, there exist constants $p_{ijL}$ with $\sup_{i,j,L}p_{ijL}^{2}<\infty$ such that
$\Psi_{\!L}=\Psi_{\!L}(K)$ $\coloneqq\sum_{i=1}^{K}\sum_{j=1}^{K}\,p_{ijL}(c_{i}\otimes d_{j}).$
Lemma A.3.
Let Assumptions 3.13.3, 3.53.6 hold. Further, suppose
$\Lambda_{K}L^{3/2}=\mathrm{o}(N^{1/2}),\quad\sum_{\ell>L}\|\psi_{\ell}\|_{\mathcal{L}}=\mathrm{O}(\lambda_{K}\Lambda^{2}_{K}L^{2}N^{-1/2}),\quad\theta_{\!N}=\mathrm{O}(\lambda_{K}\Lambda^{2}_{K}L^{5/2}N^{-1/2}).$
Then,
$\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}\big{(}\Lambda^{2}_{K}L^{5/2}N^{-1/2}\big{)}.$
Proof.
Note that $(\hat{\lambda}_{j},\hat{c}_{j})=(\hat{\lambda}_{j}(L),\hat{c}_{j}(L))_{j}$ is the eigenpair sequence of $\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}$, and $\boldsymbol{X}^{[L]}=(X^{[L]}_{k})_{k}\subset\mathcal{H}^{L}$ is the process in (3.2). By Proposition 3.1, all eigenvalues $\lambda_{j}=\lambda_{j}(L)$ of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$ are positive. Hereinafter, we write $\hat{\mathscr{D}}=\hat{\mathscr{D}}_{\!\boldsymbol{X}^{[L]},\boldsymbol{X}},$ $\mathscr{D}=\mathscr{D}_{\!\boldsymbol{X}^{[L]},\boldsymbol{X}},$ $\hat{\mathscr{C}}=\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}$ and $\mathscr{C}=\mathscr{C}_{\!\boldsymbol{X}^{[L]}}.$ The definitions of $\hat{\Psi}_{\!L}$ and $\hat{\mathscr{C}}^{\dagger}=(\hat{\mathscr{C}}+\theta_{\!N}\mathbb{I})^{-1}$, the approximate Yule–Walker equation (3.8), and $\mathscr{C}^{\ddagger}\coloneqq\mathscr{C}\mathscr{C}^{\dagger}$ yield
$\hat{\Psi}_{\!L}-\Psi_{\!L}$ $=(\hat{\mathscr{D}}-\mathscr{D})\hat{\mathscr{C}}^{\dagger}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}+\,\mathscr{D}\Big{(}\hat{\mathscr{C}}^{\dagger}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}-\,\mathscr{C}^{\dagger}\!\coprod^{c_{K}}_{c_{1}}\Big{)}+\Big{(}\sum_{\ell>L}\psi_{\ell}\mathscr{C}^{1-\ell}_{\!\!\boldsymbol{X}^{[L]},\boldsymbol{X}}\Big{)}\mathscr{C}^{\dagger}\!\coprod^{c_{K}}_{c_{1}}+\,\Psi_{\!L}\Big{(}\mathscr{C}^{\ddagger}\!\coprod^{c_{K}}_{c_{1}}-\,\mathbb{I}\Big{)}.$
The triangle inequality and the operator-valued Hölder inequality in Lemma A.1 give
$\begin{split}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}&\leq\|\hat{\mathscr{D}}-\mathscr{D}\|_{\mathcal{S}}\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}+\|\mathscr{D}\|_{\mathcal{S}}\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}-\,\mathscr{C}^{\dagger}\!\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}\\ &\qquad+\Big{\|}\sum_{\ell>L}\psi_{\ell}\mathscr{C}^{1-\ell}_{\!\!\boldsymbol{X}^{[L]},\boldsymbol{X}}\Big{\|}_{\mathcal{S}}\Big{\|}\mathscr{C}^{\dagger}\!\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}+\Big{\|}\Psi_{\!L}\Big{(}\mathscr{C}^{\ddagger}\!\coprod^{c_{K}}_{c_{1}}-\;\mathbb{I}\Big{)}\Big{\|}_{\mathcal{S}}.\end{split}$ (A.5)
According to the definition of the operator norm, for all $K,L,N$, it holds
$\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}=\Big{\|}(\hat{\mathscr{C}}+\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{-1}\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}=\sup_{j\leq K}(\hat{\lambda}_{j}+\theta_{\!N})^{-1}=(\hat{\lambda}_{K}+\theta_{\!N})^{-1},$ (A.6)
$\Big{\|}\mathscr{C}^{\dagger}\!\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}=\Big{\|}(\mathscr{C}+\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{-1}\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}=\sup_{j\leq K}(\lambda_{j}+\theta_{\!N})^{-1}=(\lambda_{K}+\theta_{\!N})^{-1}.$ (A.7)
With $K_{b}$ in the definition of $\Lambda_{K}=(\lambda_{K}-\lambda_{K_{b}})^{-1}$ in (3.11), we have
$\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}-\;\mathscr{C}^{\dagger}\!\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}$ $\leq\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\!\coprod^{\hat{c}_{K_{b}-1}}_{\hat{c}_{1}}-\;\mathscr{C}^{\dagger}\!\!\coprod^{c_{K_{b}-1}}_{c_{1}}\!\Big{\|}_{\mathcal{L}}+\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\!\coprod^{\hat{c}_{K_{b}-1}}_{\hat{c}_{K+1}}-\;\mathscr{C}^{\dagger}\!\!\coprod^{c_{K_{b}-1}}_{c_{K+1}}\!\Big{\|}_{\mathcal{L}}\,.$
Further, by the definition of the given operators and the triangle inequality,
$\begin{split}\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\!\coprod^{\hat{c}_{K_{b}-1}}_{\hat{c}_{1}}-\;\mathscr{C}^{\dagger}\!\!\coprod^{c_{K_{b}-1}}_{c_{1}}\!\Big{\|}_{\mathcal{L}}&\leq\Big{\|}\sum^{K_{b}-1}_{i=1}\,(\hat{\lambda}_{i}+\theta_{\!N})^{-1}\Big{[}(\hat{c}_{i}\otimes\hat{c}_{i})-(c_{i}\otimes c_{i})\Big{]}\Big{\|}_{\mathcal{L}}\\ &\qquad\;+\Big{\|}\sum^{K_{b}-1}_{i=1}\,\Big{[}(\hat{\lambda}_{i}+\theta_{\!N})^{-1}-(\lambda_{i}+\theta_{\!N})^{-1}\Big{]}(c_{i}\otimes c_{i})\Big{\|}_{\mathcal{L}}.\end{split}$ (A.8)
Applying Lemmas 3.1–3.2 from Reimherr (2015) to (A.8), and using $\|\cdot\|_{\mathcal{L}}\leq\|\cdot\|_{\mathcal{S}}$ along with $(\sum_{k}a_{k})^{1/2}\leq\sum_{k}a_{k}^{1/2}$ for non-negative $a_{k}$, with the reciprocal eigengaps $\hat{\Lambda}_{K}\coloneqq(\hat{\lambda}_{K}-\hat{\lambda}_{K+1})^{-1},$ which are well-defined by Assumption 3.6, it holds
$\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\!\coprod^{\hat{c}_{K_{b}-1}}_{\hat{c}_{1}}-\;\mathscr{C}^{\dagger}\!\!\coprod^{c_{K_{b}-1}}_{c_{1}}\!\Big{\|}_{\mathcal{L}}$ $\leq 2(\hat{\lambda}_{K_{b}-1}+\theta_{\!N})^{-1}\sqrt{K_{b}}\,\|\hat{\mathscr{C}}-\mathscr{C}\|_{\mathcal{L}}\big{(}\hat{\Lambda}_{K_{b}-1}+\Lambda_{K_{b}-1}+(\lambda_{K_{b}-1}+\theta_{\!N})^{-1}+1\big{)}$
$\leq 2\hat{\Lambda}_{K_{b}-1}\sqrt{K+dL}\,\|\hat{\mathscr{C}}-\mathscr{C}\|_{\mathcal{L}}\big{(}\hat{\Lambda}_{K_{b}-1}+2\Lambda_{K}+1\big{)},$ (A.9)
where $K_{b}\leq K+dL$ is given by Assumption 3.5, and $\Lambda_{K}=\Lambda_{K_{b}-1}$ by definition. Similarly,
$\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\!\coprod^{\hat{c}_{K_{b}-1}}_{\hat{c}_{K+1}}-\;\mathscr{C}^{\dagger}\!\!\coprod^{c_{K_{b}-1}}_{c_{K+1}}\!\Big{\|}_{\mathcal{L}}$ $\leq 2\hat{\Lambda}_{K_{b}-1}\sqrt{K_{b}-K}\,\|\hat{\mathscr{C}}-\mathscr{C}\|_{\mathcal{L}}\big{(}\hat{\Lambda}_{K_{b}-1}+2\Lambda_{K_{b}-1}+1\big{)}$
$\leq 2\hat{\Lambda}_{K_{b}-1}\sqrt{dL}\,\|\hat{\mathscr{C}}-\mathscr{C}\|_{\mathcal{L}}\big{(}\hat{\Lambda}_{K_{b}-1}+2\Lambda_{K}+1\big{)}.$ (A.10)
In the following, we deduce $\hat{\Lambda}_{K_{b}-1}=\mathrm{O}_{\operatorname{\mathds{P}}}(\Lambda_{K}).$ Since
$\sup_{j\in\mathbb{N}}|\hat{\lambda}_{j}-\lambda_{j}|\leq\|\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}-\mathscr{C}_{\!\boldsymbol{X}^{[L]}}\|_{\mathcal{L}}=\mathrm{O}_{\operatorname{\mathds{P}}}(L^{3/2}N^{-1/2})$
by Lemma 4.2 in Bosq (2000) together with (3.5), and $\lambda_{K_{b}-1}=\lambda_{K},$ it holds
$\hat{\lambda}_{K_{b}-1}-\hat{\lambda}_{K_{b}}$ $=\lambda_{K}-\lambda_{K_{b}}+(\hat{\lambda}_{K_{b}-1}-\lambda_{K_{b}-1})+(\lambda_{K_{b}}-\hat{\lambda}_{K_{b}})$
$=\Lambda^{-1}_{K}+\mathrm{O}_{\operatorname{\mathds{P}}}(L^{3/2}N^{-1/2}).$
Therefore, due to the assumption $\Lambda_{K}L^{3/2}=\mathrm{o}(N^{1/2}),$ in fact
$\hat{\Lambda}_{K}=\Lambda_{K}\big{[}1+\mathrm{O}_{\operatorname{\mathds{P}}}(\Lambda_{K}L^{3/2}N^{-1/2})\big{]}^{-1}=\Lambda_{K}\big{[}1+\mathrm{o}_{\operatorname{\mathds{P}}}(1)\big{]}^{-1}=\mathrm{O}_{\operatorname{\mathds{P}}}(\Lambda_{K}).$
Further, the triangle and the operator-valued Hölder inequality, and (A.2) yield for the nuclear norm $\|\cdot\|_{\mathcal{N}}\geq\|\cdot\|_{\mathcal{S}}$ of the remainder of the sum above for any $L,$
$\Big{\|}\sum_{\ell>L}\psi_{\ell}\mathscr{C}^{1-\ell}_{\!\!\boldsymbol{X}^{[L]},\boldsymbol{X}}\Big{\|}_{\mathcal{N}}\leq\sqrt{L}\operatorname{\mathds{E}}\!\|X_{0}\|^{2}\sum_{\ell>L}\|\psi_{\ell}\|_{\mathcal{L}},$ (A.11)
where the series on the right is finite by Assumption 3.2. Finally, from the definition of the H-S norm, with $\Psi_{\!L}=\Psi_{\!L}(K)$ in Assumption A.1, $\mathscr{C}^{\ddagger}=\mathscr{C}(\mathscr{C}+\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{-1},$ and the Kronecker delta $\delta_{jk},$
[…]
$\leq\theta^{2}_{\!N}(\lambda_{K}+\theta_{\!N})^{-2}\|\Psi_{\!L}(K)\|^{2}_{\mathcal{S}}\;.$ (A.12)
Thereby, Parseval’s identity yields
$\lim_{L\rightarrow\infty}\!\|\Psi_{\!L}(K)\|^{2}_{\mathcal{S}}\leq K^{2}\sup_{i,j,L}p^{2}_{ijL}<\infty.$ (A.13)
Altogether, plugging (A.6)–(A.7) and (A.1)–(A.12) in (A.5), and using $\|\mathscr{D}\|_{\mathcal{S}}\leq\sqrt{L}\operatorname{\mathds{E}}\!\|X_{0}\|^{2}$ by (A.2), $(\lambda_{K}+\theta_{\!N})^{-1}\leq\lambda^{-1}_{K}$ and $(\hat{\lambda}_{K}+\theta_{\!N})^{-1}\leq\hat{\Lambda}_{K_{b}-1},$ for each $K,L,N,$ we obtain
$\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}}$ $\leq\hat{\Lambda}_{K_{b}-1}\|\hat{\mathscr{D}}-\mathscr{D}\|_{\mathcal{S}}+2\operatorname{\mathds{E}}\!\|X_{0}\|^{2}\hat{\Lambda}_{K_{b}-1}\sqrt{(K+dL)L}\,\|\hat{\mathscr{C}}-\mathscr{C}\|_{\mathcal{L}}\,(\hat{\Lambda}_{K_{b}-1}+2\Lambda_{K}+1)$
$\quad\;\,+\lambda^{-1}_{K}\sqrt{L}\operatorname{\mathds{E}}\!\|X_{0}\|^{2}\sum_{\ell>L}\|\psi_{\ell}\|_{\mathcal{L}}+\theta_{\!N}\lambda^{-1}_{K}\|\Psi_{\!L}(K)\|_{\mathcal{S}}.$
Thus, due to $\hat{\Lambda}_{K_{b}-1}=\mathrm{O}_{\operatorname{\mathds{P}}}(\Lambda_{K}),$ Assumption 3.3, our requirements and (A.13), the claim follows. ∎
Proof of Theorem 3.1.
Follows from $\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}\leq\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}}+\|\Psi_{\!L}(K)-\Psi_{\!L}\|_{\mathcal{S}},$ the arguments to prove Lemma A.3, $\|\Psi_{\!L}(K)-\Psi_{\!L}\|_{\mathcal{S}}=\mathrm{O}(K^{-\beta})$ and $\|\Psi_{\!L}\|^{2}_{\mathcal{S}}=\sum_{i=1}^{\infty}\|\psi_{j}\|^{2}_{\mathcal{S}}<\infty$ (Remark 3.2 (c)) by Assumption 3.4, along with the requirements of our sequences. ∎
A.2 Proofs of results in Section 4
Proof of Proposition 4.1.
(a)  According to Fubini’s theorem, and as linear operators commute with the expected value, we have for any $i,j,k,$
$\operatorname{\mathds{E}}\langle c_{k},\varepsilon_{i-j}\rangle\sum_{\ell=0}^{\infty}A_{\ell}(\varepsilon_{i-\ell})$ $=\sum_{\ell=0}^{\infty}A_{\ell}(\operatorname{\mathds{E}}\langle c_{k},\varepsilon_{i-j}\rangle\varepsilon_{i-\ell}).$ (A.14)
We note that $\operatorname{\mathds{E}}\langle c_{k},\varepsilon_{i-j}\rangle\varepsilon_{i-\ell}=0$ when $j\neq\ell$, and when $j=\ell$,
$\operatorname{\mathds{E}}\langle c_{k},\varepsilon_{i-j}\rangle\varepsilon_{i-j}=\sum_{r=0}^{\infty}\operatorname{\mathds{E}}\langle c_{k},\varepsilon_{i-j}\rangle\langle c_{r},\varepsilon_{i-j}\rangle c_{r}=\lambda_{k}c_{k}$
after Bosq (2000), equation (1.36). Combining with (A.14), we have that
$\operatorname{\mathds{E}}\langle c_{k},\varepsilon_{i-j}\rangle\sum_{\ell=0}^{\infty}A_{\ell}(\varepsilon_{i-\ell})=\lambda_{k}A_{j}(c_{k}),\quad\mbox{ and }\quad\operatorname{\mathds{E}}\langle c_{k},\varepsilon_{i-j}\rangle\sum_{\ell=0}^{\infty}B_{\ell}(\varepsilon_{i-\ell})=\lambda_{k}B_{j}(c_{k}).$
Hence, due to $\lambda_{k}A_{j}(c_{k})=\lambda_{k}B_{j}(c_{k})$ for all $j,k\in\mathbb{N},$ and as $\mathscr{C}_{\boldsymbol{\varepsilon}}$ is injective, $\lambda_{k}>0$, thus $A_{j}(c_{k})=B_{j}(c_{k})$ for all $j,k\in\mathbb{N}.$ Consequently, as $(c_{j})$ forms a CONS of $\mathcal{H}$, for all $j$ holds indeed $A_{j}=B_{j}.$
(b)  Note that if $(X_{k})$ is a centered linear process, then $Z_{i}=\sum_{\ell=0}^{\infty}A_{\ell}(X_{i-\ell})$ is also one, since
$Z_{i}=\sum^{\infty}_{\ell=0}A_{\ell}(X_{i-\ell})=\sum^{\infty}_{\ell=0}D_{\ell}(\varepsilon_{i-\ell}),$
where $D_{\ell}=\sum_{j=0}^{\ell}A_{j}\phi_{\ell-j}.$ By hypothesis we also have that
$Z_{i}=\sum^{\infty}_{\ell=0}B_{\ell}(X_{i-\ell})=\sum^{\infty}_{\ell=0}F_{\ell}(\varepsilon_{i-\ell}),$
where $F_{\ell}=\sum_{j=0}^{\ell}B_{j}\phi_{\ell-j}.$ According to part (a), $F_{\ell}=D_{\ell}$ for all $\ell\in\mathbb{N}$. Since $\phi_{0}=\mathbb{I}$ we have then that $A_{0}=D_{0}=F_{0}=B_{0}$. Now the result follows by induction. ∎
We now turn towards proving Proposition 4.2. Under Assumption A.1 it follows for any $\ell$ and $K$ and $L$ sufficiently large that
$\psi_{\ell}=\psi_{\ell}(K)\coloneqq\big{[}\Psi_{\!L}(K)\big{]}^{(\ell)}=\sum_{i=1}^{K}\sum_{j=1}^{K}\,p_{ijL}(c_{i}\otimes d_{j})^{(\ell)}\,=\sum_{i=1}^{K}\sum_{j=1}^{K}\,p_{ijL}(c^{(\ell)}_{i}\!\otimes d_{j}).$ (A.15)
The finite-dimensional representations of $\phi_{i}$ need not be straightforward. For instance, for $\phi_{2}$, using $\phi_{1}=\psi_{1}$ and $\phi_{2}=\psi_{1}\phi_{1}$ under Assumption A.1 for sufficiently large $L$, we obtain
$\phi_{2}(K)=\psi^{2}_{1}(K)$ $=\sum_{i=1}^{K}\sum_{i^{\prime}=1}^{K}\sum_{j=1}^{K}\sum_{j^{\prime}=1}^{K}\,p_{ijL}\,p_{i^{\prime}j^{\prime}L}\underbrace{\big{(}c^{(\ell)}_{i}\!\otimes d_{j}\big{)}\big{(}c^{(\ell)}_{i^{\prime}}\!\otimes d_{j^{\prime}}\big{)}}_{=\;\langle c^{(\ell)}_{i}\!,\,d_{j^{\prime}}\rangle\,c^{(\ell)}_{i^{\prime}}\otimes\,d_{j}}.$
Nevertheless, we can use the recursively defined finite-dimensional approximations
$\phi_{i}(K)\coloneqq\sum^{i}_{j=1}\,\psi_{j}(K)\phi_{i-j}(K),\quad i\in\mathbb{N},\qquad\text{with}\;\,\phi_{0}(K)\coloneqq\mathbb{I}.$ (A.16)
Proposition A.1.
Let the assumptions of Lemma A.3 hold. Then,
$\|\hat{\phi}_{i}-\phi_{i}(K)\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}\big{(}\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}}\big{)}=\mathrm{O}_{\operatorname{\mathds{P}}}\big{(}\Lambda^{2}_{K}L^{5/2}N^{-1/2}\big{)},\quad i,K\in\mathbb{N}.$
Proof.
Due to the definition of $\hat{\phi_{i}}$ in (4.2) and $\phi_{i}(K)$ in (A.16), and since for each $L\geq j$ and $K$, $\|\hat{\psi_{j}}-\psi_{j}(K)\|_{\mathcal{S}}\leq\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}},$ elementary conversions give for each $i\leq L$ and $K\colon$
$\|\hat{\phi}_{i}-\phi_{i}(K)\|_{\mathcal{S}}$ $\leq\sum^{i-1}_{j=1}\,\|\hat{\psi}_{j}\|_{\mathcal{L}}\|\hat{\phi}_{i-j}-\phi_{i-j}(K)\|_{\mathcal{S}}+\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}}\sum^{i-1}_{j=0}\,\|\phi_{j}\|_{\mathcal{L}},$
where we used $\hat{\phi}_{0}=\phi_{0}(K)=\mathbb{I}$ and $\|\phi_{j}(K)\|_{\mathcal{L}}\leq\|\phi_{j}\|_{\mathcal{L}}$. Since $\|\hat{\phi}_{1}-\phi_{1}(K)\|_{\mathcal{S}}=\|\hat{\psi}_{1}-\psi_{1}(K)\|_{\mathcal{S}}$ is bounded above by $\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}}$, whose asymptotic upper bound follows from Lemma A.3, iterating the above $i-1$ times yields an asymptotic bound for $\|\hat{\phi}_{i}-\phi_{i}(K)\|_{\mathcal{S}}$. Namely, with
$\Theta(i)\coloneqq 1+\sum^{i-1}_{k=1}\sum^{i-j_{0}-1}_{j_{1}=1}\cdots\sum^{i-j_{k-1}-1}_{j_{k}=1}\!\Big{(}\prod^{k-1}_{\ell=1}\,\|\psi_{j_{k}}\|_{\mathcal{L}}\!\Big{)}\Big{[}\,\|\phi_{j_{k}}\|_{\mathcal{L}}+\|\psi_{j_{k}}\|_{\mathcal{L}}\Big{]},\quad i\in\mathbb{N},$ (A.17)
and $\|\phi_{0}\|_{\mathcal{L}}=1,$ $\|\hat{\psi}_{j}\|_{\mathcal{L}}=\|\psi_{j}(K)\|_{\mathcal{L}}+\mathrm{o}_{\operatorname{\mathds{P}}}(1)$ for all $j,K$ by Lemma A.3 and $\|\psi_{j}(K)\|_{\mathcal{L}}\leq\|\psi_{j}\|_{\mathcal{L}},$
$\|\hat{\phi}_{i}-\phi_{i}(K)\|_{\mathcal{S}}$ $\leq\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}}\bigg{[}\sum^{i^{\prime}}_{j_{1}=0}\|\phi_{j_{1}}\|_{\mathcal{L}}+\sum^{i^{\prime}}_{j_{1}=1}\sum^{i^{\prime}-j_{1}}_{j_{2}=0}\|\hat{\psi}_{j_{1}}\|_{\mathcal{L}}\|\phi_{j_{2}}\|_{\mathcal{L}}\bigg{.}$
.$\bigg{.}\qquad\qquad\qquad\qquad\quad\;\;+\cdots+\sum^{i^{\prime}}_{j_{1}=1}\sum^{i^{\prime}-j_{1}}_{j_{2}=1}\cdots\sum^{i^{\prime}-j_{i-3}}_{j_{i-2}=1}\sum^{i^{\prime}-j_{i-2}}_{j_{i-1}=0}\!\Big{(}\prod^{i^{\prime}-1}_{k=1}\|\hat{\psi}_{j_{k}}\|_{\mathcal{L}}\Big{)}\|\phi_{j_{i-1}}\|_{\mathcal{L}}\bigg{]}\allowdisplaybreaks$
$\qquad\qquad+\sum^{i^{\prime}}_{j_{1}=1}\sum^{i^{\prime}-j_{1}}_{j_{2}=1}\cdots\sum^{i^{\prime}-j_{i-2}}_{j_{i-1}=1}\!\Big{(}\prod^{i^{\prime}}_{k=1}\|\hat{\psi}_{j_{k}}\|_{\mathcal{L}}\Big{)}\underbrace{\;\big{\|}\hat{\phi}_{i-\sum^{i-1}_{k=1}j_{k}}\!-\phi_{i-\sum^{i-1}_{k=1}j_{k}}\big{\|}_{\mathcal{S}}\;}_{=\;\|\hat{\psi}_{1}-\psi_{1}(K)\|_{\mathcal{S}}}\allowdisplaybreaks$
$\leq\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}}\big{[}\Theta(i)+\mathrm{o}_{\operatorname{\mathds{P}}}(1)\big{]},$ (A.18)
with $i^{\prime}\coloneqq i-1$, where factors in multi-sums with $\sum^{0}_{j_{k}=1}$ for some $k$ were set to 1. ∎
Proof of Proposition 4.2.
Similar to the arguments in the proof of Proposition A.1, and since $\|\psi_{j}(K)-\psi_{j}\|_{\mathcal{S}}\leq\|\Psi_{\!L}(K)-\Psi_{\!L}\|_{\mathcal{S}}$ for all $j,K,$ for $i\leq L,$ it holds that
$\|\phi_{i}(K)-\phi_{i}\|_{\mathcal{S}}$ $\leq\sum^{i-1}_{j=1}\,\|\psi_{j}\|_{\mathcal{L}}\|\phi_{i-j}(K)-\phi_{i-j}\|_{\mathcal{S}}+\|\Psi_{\!L}(K)-\Psi_{\!L}\|_{\mathcal{S}}\sum^{i-1}_{j=0}\,\|\phi_{j}\|_{\mathcal{L}}.$
Performing $i-1$ iterations of this inequality, following the rationale to prove Proposition A.1, and using $\|\phi_{1}(K)-\phi_{1}\|_{\mathcal{S}}=\|\psi_{1}(K)-\psi_{1}\|_{\mathcal{S}},$ gives $\|\phi_{i}(K)-\phi_{i}\|_{\mathcal{S}}\leq\|\Psi_{\!L}(K)-\Psi_{\!L}\|_{\mathcal{S}}\Theta(i),$ with $\Theta(i)$ in (A.17). Further, (A.18), $\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta})$ and $\|\Psi_{\!L}(K)-\Psi_{\!L}\|_{\mathcal{S}}=\mathrm{O}(K^{-\beta})$ due to the requirements of Theorem 3.1, and $\|\hat{\phi}_{i}-\phi_{i}\|_{\mathcal{S}}\leq\|\hat{\phi}_{i}-\phi_{i}(K)\|_{\mathcal{S}}+\|\phi_{i}(K)-\phi_{i}\|_{\mathcal{S}}$ prove our claim. ∎
A.3 Proofs of results in Section 5
Proof of Corollary 5.1.
Follows directly from Theorem 3.1 with $L=p.$
Proof of Corollary 5.2.
Follows from Proposition 4.2 by putting $\beta_{j}=\phi_{j}$ and $\hat{\beta}_{j}=\hat{\phi}_{j}.$
To derive the asymptotic result for our fARMA(1,1)11(1,1)( 1 , 1 ) operators, we derive a preliminary result based on the following assumption.
Assumption A.2.
For fixed $M\in\mathbb{N},$ there are constants $b_{ij}\in\mathbb{R}$ so that
$\beta_{1}=\beta_{1}(M)$ $\coloneqq\sum_{i=1}^{M}\sum_{j=1}^{M}\,b_{ij}(f_{i}\otimes f_{j}).$
A finite-dimensional representation of $\beta_{1}$ does not imply one for $\alpha_{1}$ under Assumption 5.1. As (5.3) gives $\alpha_{1}=\alpha_{1}(M)\coloneqq\psi_{1}+\beta_{1}(M)$, assuming $\psi_{1}$ is finite-dimensional contradicts Assumption 5.1.
Lemma A.4.
Suppose $(X_{k})$ is the fARMA(1,1)11(1,1)( 1 , 1 ) process in (5.2). Further, let the conditions of Theorem 3.1, Assumptions 5.1, 5.35.4, A.2, and $\gamma_{N}=\mathrm{O}(K^{-\beta})$ hold. Then,
$\max\!\big{\{}\|\hat{\alpha}_{1}-\alpha_{1}(M)\|_{\mathcal{S}},\|\hat{\beta}_{1}-\beta_{1}(M)\|_{\mathcal{S}}\big{\}}$ $=\mathrm{O}_{\operatorname{\mathds{P}}}\big{(}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}\big{)}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta}).$
Proof.
The conversions here follow those in Lemma A.3. Recall $(\gamma_{N})\subset(0,\infty)$ with $\gamma_{N}\to 0$ and $\beta_{1}=\beta_{1}(M)$ by Assumption A.2. Let $(\rho_{j},f_{j})$ and $(\hat{\rho}_{j},\hat{f}_{j})$ be the eigenpair sequences of $\psi_{1}\psi_{1}^{\ast}$ and $\hat{\psi}_{1}\hat{\psi}_{1}^{\ast}$, respectively, where $\psi_{i}$ are the H-S operators from (3.1), and $\hat{\psi}_{1}=\hat{\Psi}^{(i)}_{\!L}$ its estimates from (3.10). Since $\psi_{1}\psi_{1}^{\ast}$ is injective (Assumption 5.1), all $\rho_{j}$ are positive. Moreover,
$\|\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}-\psi_{1}\psi^{\ast}_{1}\|_{\mathcal{L}}\leq\|\hat{\psi}_{1}-\psi_{1}\|_{\mathcal{L}}\big{(}\|\psi_{1}\|_{\mathcal{L}}+\|\hat{\psi}_{1}\|_{\mathcal{L}}\big{)}.$
As $\hat{\beta}_{1}\coloneqq-\hat{\psi}_{2}\hat{\psi}^{\dagger}_{1}\coprod^{\hat{f}_{M}}_{\hat{f}_{1}},$ where $\hat{\psi}^{\dagger}_{1}=\hat{\psi}^{\ast}_{1}(\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}+\gamma_{N}\mathbb{I})^{-1}\!,$ with $\psi^{\dagger}_{1}\coloneqq\psi^{\ast}_{1}(\psi_{1}\psi^{\ast}_{1}+\gamma_{N}\mathbb{I})^{-1}\!,\psi^{\ddagger}_{1}\coloneqq\psi^{\ast}_{1}\psi^{\dagger}_{1},$ and $\psi_{2}=-\beta_{1}\psi_{1}$ from (5.3), due to triangle and operator-valued Hölder’s inequality, it holds
$\|\hat{\beta}_{1}-\beta_{1}\|_{\mathcal{S}}\leq\|\hat{\psi}_{2}-\psi_{2}\|_{\mathcal{S}}\Big{\|}\hat{\psi}^{\dagger}_{1}\coprod^{\hat{f}_{M}}_{\hat{f}_{1}}\Big{\|}_{\mathcal{L}}+\|\psi_{2}\|_{\mathcal{S}}\Big{\|}\hat{\psi}^{\dagger}_{1}\coprod^{\hat{f}_{M}}_{\hat{f}_{1}}-\psi^{\dagger}_{1}\coprod^{f_{M}}_{f_{1}}\Big{\|}_{\mathcal{L}}+\Big{\|}\beta_{1}\Big{(}\psi^{\ddagger}_{1}\coprod^{f_{M}}_{f_{1}}-\mathbb{I}\Big{)}\Big{\|}_{\mathcal{S}}\;.$
Due to the properties of our operators and their norms, it follows that
$\Big{\|}\hat{\psi}^{\dagger}_{1}\coprod^{\hat{f}_{M}}_{\hat{f}_{1}}\Big{\|}^{2}_{\mathcal{L}}$ $=\Big{\|}\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}\big{(}\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}+\gamma_{N}\mathbb{I}\big{)}^{-2}\coprod^{\hat{f}_{M}}_{\hat{f}_{1}}\Big{\|}_{\mathcal{L}}\leq(\hat{\rho}_{M}+\gamma_{N})^{-1}.$
Further, from the definition of $\hat{\psi}^{\dagger}_{1}$ and $\psi^{\dagger}_{1},$ since $\|A^{\ast}\|_{\mathcal{L}}=\|A\|_{\mathcal{L}}$ for bounded linear operators $A,$ elementary transformations, and similar arguments as in the proof of Lemma A.3, it follows
$\Big{\|}\hat{\psi}^{\dagger}_{1}\coprod^{\hat{f}_{M}}_{\hat{f}_{1}}-\,\psi^{\dagger}_{1}\coprod^{f_{M}}_{f_{1}}\Big{\|}_{\mathcal{L}}$
$\;\leq\|\hat{\psi}_{1}\|_{\mathcal{L}}\Big{\|}(\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}+\gamma_{N}\mathbb{I})^{-1}\coprod^{\hat{f}_{M}}_{\hat{f}_{1}}-\,(\psi_{1}\psi^{\ast}_{1}+\gamma_{N}\mathbb{I})^{-1}\coprod^{f_{M}}_{f_{1}}\Big{\|}_{\mathcal{L}}+\|\hat{\psi}_{1}-\psi_{1}\|_{\mathcal{L}}\Big{\|}(\psi_{1}\psi^{\ast}_{1}+\gamma_{N}\mathbb{I})^{-1}\coprod^{f_{M}}_{f_{1}}\Big{\|}_{\mathcal{L}}$
$\;\leq 2\hat{P}_{M_{b}-1}\sqrt{M+d}\,\|\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}-\psi_{1}\psi^{\ast}_{1}\|_{\mathcal{L}}\big{(}\hat{P}_{M_{b}-1}+2P_{M}+1\big{)}+\rho^{-1}_{\!M}\|\hat{\psi}_{1}-\psi_{1}\|_{\mathcal{L}},$
with $\hat{P}_{M_{b}-1}\coloneqq(\hat{\rho}_{M_{b}-1}-\hat{\rho}_{M_{b}})^{-1},$ and $P_{M}\coloneqq(\rho_{M}-\rho_{M_{b}})^{-1}$ as in (3.11), where former is well-defined by Assumption 5.4. We also used that $M_{b}\leq M+d$ for some $d\in\mathbb{N}$ by Assumption 5.3. Moreover, as in the proof of Lemma A.3, with $\beta_{1}=\beta_{1}(M)$ in Assumption A.2, it holds
$\Big{\|}\beta_{1}\Big{(}\psi^{\ddagger}_{1}\coprod^{f_{M}}_{f_{1}}-\mathbb{I}\Big{)}\Big{\|}_{\mathcal{S}}$ $\leq\gamma_{N}\rho^{-1}_{\!M}\|\beta_{1}(M)\|_{\mathcal{S}}.$
Altogether, with $\|\cdot\|_{\mathcal{L}}\leq\|\cdot\|_{\mathcal{S}},$ $\|\hat{\psi}_{i}-\psi_{i}\|_{\mathcal{S}}\leq\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}$ for all $i$, and $(\hat{\rho}_{M}+\gamma_{N})^{-1}\leq\hat{P}_{M_{b}-1},$ $\hat{P}_{M_{b}-1}=\mathrm{O}_{\operatorname{\mathds{P}}}(P_{M})$ (see the proof of Lemma A.3), $P^{1/2}_{M}=\mathrm{O}(P_{M}),$ and $\rho^{-1}_{M}\leq P_{M},$ it holds
$\|\hat{\beta}_{1}-\beta_{1}(M)\|_{\mathcal{S}}$ $\leq\mathrm{O}_{\operatorname{\mathds{P}}}\big{(}M^{1/2}P^{2}_{M}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}\big{)}+\mathrm{O}(\gamma_{N}\rho^{-1}_{\!M}).$ (A.19)
Thus, for fixed $M,$ $\|\hat{\beta}_{1}-\beta_{1}(M)\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta})$ by Theorem 3.1 and $\gamma_{N}=\mathrm{O}(K^{-\beta}).$ Consequently, due to $\hat{\alpha}_{1}=\hat{\psi}_{1}-\hat{\beta}_{1}$ and $\alpha_{1}(M)=\psi_{1}-\beta_{1}(M),$ for fixed $M,$ we also obtain $\|\hat{\alpha}_{1}-\alpha_{1}(M)\|_{\mathcal{S}}\!\leq\!\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}+\|\hat{\beta}_{1}-\beta_{1}(M)\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{\-\beta}).$
Proof of Theorem 5.1.
Due to (A.19), $M^{\gamma+1/2}P^{2}_{M}=\mathrm{O}(K^{\beta})$, $\gamma_{N}=\mathrm{O}(\rho_{M}M^{-\gamma})$ and Assumption 5.2, we have $\|\hat{\beta}_{1}-\beta_{1}\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}(M^{-\gamma})$ (see the proof of Theorem 3.1). In consequence, due to $\alpha_{1}=\psi_{1}+\beta_{1},$ Theorem 3.1, and $K^{-\beta}=\mathrm{O}(M^{-\gamma}),$ it also holds $\|\hat{\alpha}_{1}-\alpha_{1}\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}(M^{-\gamma})$. ∎
Assumption A.3.
For fixed $M\in\mathbb{N},$ there are constants $b_{ijq}\in\mathbb{R}$ so that
$B_{q}=B_{q}(M)$ $\coloneqq\sum_{i=1}^{M}\sum_{j=1}^{M}\,b_{ijq}(h_{i}\otimes d_{j}).$
From (5.5), which is equivalent to $\alpha_{i}=\psi_{i}+B_{[i]}\Psi^{\prime\prime\prime\,\top}_{[i]}$ for $i=1,\dots,p,$ as in Section 5.3, it follows
$\alpha_{i}=\alpha_{i}(M)\coloneqq\psi_{i}+B_{[i]}(M)\Psi^{\prime\prime\,\top}_{[i]},\quad 1\leq i\leq p.$
Lemma A.5.
Let $(X_{k})$ be the fARMA$(p,q)$ process in (5.1). Moreover, let the conditions of Theorem 3.1, Assumptions 5.5, 5.75.8, A.3 and $\gamma_{N}=\mathrm{O}(K^{-\beta})$ hold. Then,
$\max_{1\leq i\leq p,1\leq j\leq q}\!\big{\{}\|\hat{\alpha}_{i}-\alpha_{i}(M)\|_{\mathcal{S}},\|\hat{\beta}_{j}-\beta_{j}(M)\|_{\mathcal{S}}\big{\}}=\mathrm{O}_{\operatorname{\mathds{P}}}\big{(}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}\big{)}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta}).$
Proof.
The proof follows that of Lemma A.4. Let $(\zeta_{j},h_{j})$ and $(\hat{\zeta}_{j},\hat{h}_{j})$ be the eigenpair sequences of ,^^$\boldsymbol{\prod}\boldsymbol{\prod}^{\ast},\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}\in\mathcal{N}_{\mathcal{H}^{q}},$ with ,^$\boldsymbol{\prod},\boldsymbol{\hat{\prod}}\in\mathcal{S}_{\mathcal{H}^{q}}$ in (5.8) and (5.10), respectively. Under Assumption 3.4 and Lemma A.2 (b), we have $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}\in\mathcal{N}_{\mathcal{H}^{q}},$ and Assumption 5.5 ensures all eigenvalues $\zeta_{j}$ are positive. Finally, basic conversions and the definitions of our operator-valued vectors and matrices yield
$\Big{\|}\boldsymbol{\hat{\prod}}-\boldsymbol{\prod}\Big{\|}_{\mathcal{L}}$ $\leq\sum^{q-1}_{i=0}\,\|\hat{\Psi}^{\prime}_{[i]}-\Psi^{\prime}_{[i]}\|_{\mathcal{L}}\,.$
Consequently, as $\|\hat{\Psi}^{\prime}_{[i]}-\Psi^{\prime}_{[i]}\|_{\mathcal{S}}\leq\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}$ for $L\geq p+q+i-1,$ we obtain for $L\geq p+2q-2,$
$\Big{\|}\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}\!-\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}\Big{\|}_{\mathcal{L}}$ $\leq q\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}\Big{(}\Big{\|}\boldsymbol{\prod}\Big{\|}_{\mathcal{L}}+\Big{\|}\boldsymbol{\hat{\prod}}\Big{\|}_{\mathcal{L}}\Big{)}.$ (A.20)
With $\hat{B}_{q}=-\hat{\Psi}^{\prime}_{[q]}\boldsymbol{\hat{\prod}}^{\dagger}\!\coprod^{\hat{h}_{\!M}}_{\hat{h}_{1}},$ where ^=^(^^$\boldsymbol{\hat{\prod}}^{\dagger}=\boldsymbol{\hat{\prod}}^{\ast}(\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}+\gamma_{N}\mathbb{I})^{-1}$ and $\boldsymbol{\prod}^{\dagger}=\boldsymbol{\prod}^{\ast}(\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}+\gamma_{N}\mathbb{I})^{-1},$ with $(\gamma_{N})\subset(0,\infty)$ satisfying $\gamma_{N}\to 0,$ and with $\boldsymbol{\prod}^{\ddagger}\coloneqq\boldsymbol{\prod}\boldsymbol{\prod}^{\dagger}$ and $\Psi^{\prime}_{[q]}=-B_{q}\boldsymbol{\prod}$ in (5.7), for $B_{q}=B_{q}(M)$ in Assumption A.3, we obtain
$\hat{B}_{q}-B_{q}(M)$ $=\big{(}\Psi^{\prime}_{[q]}-\hat{\Psi}^{\prime}_{[q]}\big{)}\boldsymbol{\hat{\prod}}^{\dagger}\coprod^{\hat{h}_{\!M}}_{\hat{h}_{1}}+\,\Psi^{\prime}_{[q]}\Big{(}\boldsymbol{\prod}^{\dagger}\!\coprod^{h_{\!M}}_{h_{1}}-\boldsymbol{\hat{\prod}}^{\dagger}\!\coprod^{\hat{h}_{\!M}}_{\hat{h}_{1}}\Big{)}+B_{q}\Big{(}\boldsymbol{\prod}^{\ddagger}\!\coprod^{h_{\!M}}_{h_{1}}-\,\mathbb{I}\Big{)}.$
Further, as in the proof of Lemma A.4, together with (A.20) and Assumptions 5.75.8, it holds
$\|\hat{B}_{q}-B_{q}(M)\|_{\mathcal{S}}$ $\leq(\hat{\zeta}_{M}+\gamma_{N})^{-1/2}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}+\|\Psi^{\prime}_{[q]}\|_{\mathcal{S}}\Big{\|}\boldsymbol{\prod}^{\dagger}\!\coprod^{h_{\!M}}_{h_{1}}-\boldsymbol{\hat{\prod}}^{\dagger}\!\coprod^{\hat{h}_{\!M}}_{\hat{h}_{1}}\Big{\|}_{\mathcal{L}}+\gamma_{N}\zeta^{-1}_{M}\|B_{q}(M)\|_{\mathcal{S}}$
$\leq\mathrm{O}_{\operatorname{\mathds{P}}}\big{(}M^{1/2}Z^{2}_{M}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}\big{)}+\mathrm{O}(\gamma_{N}\zeta^{-1}_{\!M})$
Hence, as $\gamma_{N}=\mathrm{O}(K^{-\beta}),$ for fixed $M$ holds $\|\hat{B}_{q}-B_{q}(M)\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta})$ and thus $\|\hat{\beta}_{j}-\beta_{j}(M)\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta})$ for all $j.$ Further, $\hat{\alpha}_{i}=\hat{\psi}_{i}+\hat{B}_{[i]}\hat{\Psi}^{\prime\prime\prime\,\top}_{[i]},$ $\alpha_{i}(M)=\psi_{i}+B_{[i]}(M)\Psi^{\prime\prime\prime\,\top}_{[i]}\!,$ and $\alpha_{i}=\beta_{j}=\hat{\beta}_{j}=0$ for $i>p,$ $j>q,$ give for $L\geq p+q+i-1$ for any $i,p,q\colon$
$\|\hat{\alpha}_{i}-\alpha_{i}(M)\|_{\mathcal{S}}$ $\leq\|\hat{\psi}_{i}-\psi_{i}\|_{\mathcal{S}}+\|\hat{B}_{[i]}(M)\|_{\mathcal{L}}\|\hat{\Psi}^{\prime\prime\prime}_{[i]}-\Psi^{\prime\prime\prime}_{[i]}\|_{\mathcal{S}}+\|\Psi^{\prime\prime\prime}_{[i]}\|_{\mathcal{L}}\|\hat{B}_{[i]}-B_{[i]}(M)\|_{\mathcal{S}}$
$\leq(1+\|\hat{B}_{[i]}(M)\|_{\mathcal{L}})\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}+\|\Psi^{\prime\prime\prime}_{[i]}\|_{\mathcal{L}}\|\hat{B}_{q}-B_{q}(M)\|_{\mathcal{S}}.$
$\leq\|\hat{B}_{q}-B_{q}(M)\|_{\mathcal{S}}\big{(}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}+\|\Psi^{\prime\prime\prime}_{[i]}\|_{\mathcal{L}}\big{)}+\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}},$
and consequently also $\|\hat{\alpha}_{i}-\alpha_{i}(M)\|_{\mathcal{S}}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta})$ for all $i.$ This shows our claim. ∎
Proof of Theorem 5.2.
Follows from the argumentation in the proofs of Theorem 5.1 and Lemma A.5 and the requirements for our sequences. ∎

### v3

@@
 These results further provide consistency-rate estimates for operators in Hilbert space-valued causal linear processes, including functional MA, AR, and ARMA models of arbitrary order.
-MSC 2020 subject classifications: 47B38, 60G10, 62F12
+MSC 2020 subject classifications: 60G10, 62F12, 62R10
 Keywords: ARMA; functional time series; invertible processes; linear processes
@@
 Turbillon et al.
-(2008) estimated the fMA$(1)$ operator assuming commutativity, and Kuenzer (2024) derived consistent estimators for fARMA operators using a two-step approach, involving sequential AR and MA estimation with dimension reduction via principal components.
+(2008) estimated the fMA(1)(1)( 1 ) operator assuming commutativity, and Kuenzer (2024) derived consistent estimators for fARMA operators using a two-step approach, involving sequential AR and MA estimation with dimension reduction via principal components.
 Central limit theorems for fAR operator estimates are derived in Mas and Pumo (2009).
@@
 Notably, our results are derived under mild conditions, permitting multi-dimensional eigenspaces of the covariance operator, and without requiring specific rates of eigenvalue decay.
-Provided the operators exhibit a sufficiently high degree of smoothness and the eigenvalue decay of the error covariance operators is slow, our approach yields the near-parametric rate $N^{-1/2}.$
+Provided the operators exhibit a sufficiently high degree of smoothness and the eigenvalue decay of the error covariance operators is slow, our approach nearly achieves the parametric rate $N^{-1/2}.$
 The paper is organized as follows: Section 2 introduces notation and fundamental concepts.
@@
 2.1 Notation
-The additive identity in linear spaces is denoted by 00, and the identity map by $\mathbb{I}$.
+The additive identity in linear spaces is denoted by 0, and the identity map by $\mathbb{I}$.
 In Cartesian product spaces $V^{n}$, $n\in\mathbb{N}$, scalar multiplication and addition are defined component-wise.
@@
 Assumption 3.2.
-$(X_{k})_{k\in\mathbb{Z}}\subset L^{2}_{\mathcal{H}}$ is a centered, invertible process with representation (3.1), and the operators $\psi_{j}\in\mathcal{S}_{\mathcal{H}}$.
+$(X_{k})_{k\in\mathbb{Z}}\subset L^{2}_{\mathcal{H}}$ is a centered, invertible process with representation (3.1), with operators $\psi_{j}\in\mathcal{S}_{\mathcal{H}}$.
 Consistent estimates for the operators $\psi_{j}$ are derived based on a sample $X_{1},\dots,X_{N}$ from ${\boldsymbol{X}}=(X_{k})$.
@@
 $X_{k}=\sum_{\ell=-\infty}^{\infty}\phi_{\ell}(\varepsilon_{k-\ell}),$ (3.7)
-where ϕsubscriptitalic-$\phi_{\ell}\in\mathcal{L_{H}}$ with $\sum^{\infty}_{\ell=-\infty}\|\phi_{\ell}\|_{\mathcal{L}}<\infty.$ Then, Assumption 3.3 is satisfied if the linear process operators fulfill the stronger condition $\sum^{\infty}_{k=1}\sum^{\infty}_{\ell=k}\left[\|\phi_{\ell}\|_{\mathcal{L}}+\|\phi_{-\ell}\|_{\mathcal{L}}\right]<\infty.$
+where $\phi_{\ell}\in\mathcal{L_{H}}$ with $\sum^{\infty}_{\ell=-\infty}\|\phi_{\ell}\|_{\mathcal{L}}<\infty.$ Then, Assumption 3.3 is satisfied if the linear process operators fulfill the stronger condition $\sum^{\infty}_{k=-\infty}|k|\,\|\phi_{k}\|_{\mathcal{L}}<\infty.$
 The estimates for the operators $\psi_{j}$ in the inverted representation (3.1) are derived based on the approximate Yule–Walker equation
@@
 For some $\beta>0,$ the limit $S_{\Psi}(\beta)\coloneqq\lim_{L\rightarrow\infty}S_{\Psi_{\!L}}(\beta)<\infty$ exists, where for each $L=L_{N}\in\mathbb{N}$,
-$S_{\Psi_{\!L}}\!(\beta)\coloneqq\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\Psi_{\!L},c_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}\,(1+i^{2\beta}\!+j^{2\beta})<\infty.$
+$S_{\Psi_{\!L}}\!(\beta)\coloneqq\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\Psi_{\!L},c_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}\,(1+i^{2\beta}\!+\hskip-0.50003ptj^{2\beta})<\infty.$
 Our results bellow allow for repeated eigenvalues $\lambda_{j}=\lambda_{j}(L)$ of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}},$ provided the dimension of each associated eigenspace $\mathcal{V}_{j}(L)$ of $\mathscr{C}_{\!\boldsymbol{X}^{[L]}}$ is bounded in the following sense.
@@
 Assumption 3.6.
-For all $N$ sufficiently large, with probability one the eigenvalues $\hat{\lambda}_{j}=\hat{\lambda}_{j}(L)$ of $\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}$ satisfy $\hat{\lambda}_{j}\neq\hat{\lambda}_{j+1}$ for all $j\in\{1,...,K\}$ and $L.$
+For all $N$ sufficiently large, with probability one the eigenvalues $\hat{\lambda}_{j}=\hat{\lambda}_{j}(L)$ of $\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}$ satisfy $\hat{\lambda}_{j}\neq\hat{\lambda}_{j+1}$ for all $j\in\{1,...,K+dL\}$ and $L,$ with $d$ from Assumption 3.5.
 Theorem 3.1.
@@
 5 Consequences for fARMA processes
-The asymptotic results in Section 3 extend to the estimation of fAR, fMA, and fARMA parameters.
-These models are extensively studied in Bosq (2000), particularly the fAR case, while Spangenberg (2013) provides general conditions for causal and invertible representations in Banach spaces.
-Estimation results for the fAR(1) operator are available in Bosq (2000) and Caponera and Panaretos (2022).
-Additionally, Turbillon et al.
-(2008) estimated the fMA(1) operator assuming commutativity with the error covariance operator, and Aue and Klepsch (2017); Klepsch et al.
-(2017) studied prediction for fARMA$(p,q)$ processes.
-Kuenzer (2024) derived consistent fMA(1) and fARMA(1,1) operator estimates using an adaptation of the Hannan–Rissanen method (Hannan and Rissanen, 1982).
-By using operator estimates from both the linear and inverted representations, we derive consistent fAR, fMA, and fARMA operator estimates with explicit rates.
+[…]
 5.1 fAR processes
-A centered process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a functional autoregressive process of order $p\in\mathbb{N}$ (fAR$(p)$) if
+[…]
 $X_{k}=\sum^{p}_{i=1}\,\alpha_{i}(X_{k-i})+\varepsilon_{k},\quad k\in\mathbb{Z},$
@@
 5.2 fMA processes
-A centered process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a functional moving-average process of order $q\in\mathbb{N}$ (fMA$(q)$) if
+[…]
 $X_{k}=\sum^{q}_{j=1}\,\beta_{j}(\varepsilon_{k-j})+\varepsilon_{k},\quad k\in\mathbb{Z},$
@@
 5.3 fARMA processes
-We finally turn to the estimation of fARMA processes.
-A centered process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is called a functional autoregressive moving-average process of order $p,q\in\mathbb{N}$ (fARMA$(p,q)$) if
+[…]
 $X_{k}=\sum^{p}_{i=1}\,\alpha_{i}(X_{k-i})+\sum^{q}_{j=1}\,\beta_{j}(\varepsilon_{k-j})+\varepsilon_{k},\quad k\in\mathbb{Z},$ (5.1)
-where $(\varepsilon_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a WN, and $\alpha_{i},\beta_{j}\in\mathcal{L_{H}}$ are operators with $\alpha_{p}\neq 0,\beta_{q}\neq 0.$ We combine this fARMA$(p,q)$ representation with its inverted form (3.1).
-As with multivariate ARMA processes, the operators $\alpha_{i}$ and $\beta_{j}$ are generally not identifiable.
-Sufficient conditions for identifiability are provided in Kuenzer (2024) (equation (10) and Propositions 1–3).
-We assume that the fARMA models considered are identifiable and admit invertible, centered, linear process solutions.
-To describe the main idea, we first discuss estimation for fARMA(1,1)11(1,1)( 1 , 1 ) processes of the form
+[…]
+To describe the main idea, we first discuss estimation for fARMA(1,1)(1,1)( 1 , 1 ) processes of the form
 $X_{k}=\alpha_{1}(X_{k-1})+\,\beta_{1}(\varepsilon_{k-1})+\varepsilon_{k},\quad k\in\mathbb{Z}.$ (5.2)
@@
 $\psi_{i}=\begin{cases}\,\alpha_{1}+\beta_{1},&\textnormal{if}\;\,i=1,\\ \,-\beta_{1}\psi_{i-1},&\textnormal{if}\;\,i>1.\\ \end{cases}$ (5.3)
-Although $\hat{\Psi}_{\!L}$ in (3.10) provides consistent asymptotic estimates for all $\psi_{i}$, consistent estimates for the fARMA(1,1)11(1,1)( 1 , 1 ) operators cannot be directly obtained, as $\psi_{1}\in\mathcal{S_{H}}$ does not possess a bounded inverse.
+Although $\hat{\Psi}_{\!L}$ in (3.10) provides consistent asymptotic estimates for all $\psi_{i}$, consistent estimates for the fARMA(1,1)(1,1)( 1 , 1 ) operators cannot be directly obtained, as $\psi_{1}\in\mathcal{S_{H}}$ does not possess a bounded inverse.
 However, using $\hat{\psi}_{i}=\hat{\Psi}^{(i)}_{\!L}$, equation (5.3) leads to the following estimators for $\alpha_{1}$ and $\beta_{1}$:
@@
 The operators $\alpha_{1},\beta_{1}\in\mathcal{S_{H}}$ are H-S, and for some $\gamma>0,$ it holds
-$\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\beta_{1},f_{i}\otimes f_{j}\rangle^{2}_{\mathcal{S}}(1+i^{2\gamma}\!+j^{2\gamma})<\infty.$
+$\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\beta_{1},f_{i}\otimes f_{j}\rangle^{2}_{\mathcal{S}}(1+i^{2\gamma}\!+\hskip-0.50003ptj^{2\gamma})<\infty.$
 Assumption 5.3.
-The dimensions of all eigenspaces of $\psi_{1}\psi^{\ast}_{1}$ are bounded above by some $d\in\mathbb{N}.$
-Assumption 5.4.
-For all $N$ sufficiently large, with probability one the eigenvalues of $\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}$ satisfy $\hat{\rho}_{j}\neq\hat{\rho}_{j+1}$ for all $j\in\{1,...,M\}.$
+There exists $d\in\mathbb{N}$ such that:
+  • (a)
+    The dimensions of all eigenspaces of $\psi_{1}\psi^{\ast}_{1}$ are bounded above by $d;$
+  • (b)
+    For all $N$ sufficiently large, with probability one the eigenvalues of $\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}$ satisfy $\hat{\rho}_{j}\neq\hat{\rho}_{j+1}$ for all $j\in\{1,...,M+d\}.$
 Theorem 5.1.
-Suppose $(X_{k})$ is the fARMA(1,1)11(1,1)( 1 , 1 ) process in (5.2).
+Suppose $(X_{k})$ is the fARMA(1,1)(1,1)( 1 , 1 ) process in (5.2).
 Further, let the conditions of Theorem 3.1 and Assumptions 5.15.3 hold, and let $M=M_{\!N}\rightarrow\infty$ with $M^{\gamma+1/2}P^{2}_{\!M}=\mathrm{O}(K^{\beta})$ and $\gamma_{N}=\mathrm{O}(\rho_{M}M^{-\gamma}),$ with $P_{M}\coloneqq(\rho_{M}-\rho_{M_{b}})^{-1}$ as in (3.11).
@@
 $\hat{B}_{q}\coloneqq-\hat{\Psi}^{\prime}_{[q]}\,\boldsymbol{\hat{\prod}}^{\dagger}\coprod^{\hat{h}_{\!M}}_{\hat{h}_{1}},$
-where ^(^^$\boldsymbol{\hat{\prod}}^{\dagger}\coloneqq(\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}+\gamma_{N}\hskip 0.50003pt\mathbb{I})^{-1},$ $(M_{N})\subset\mathbb{N}$ and $(\gamma_{N})\subset(0,\infty)$ are sequences with $M_{N}\to\infty$ and $\gamma_{N}\rightarrow 0,$ and $\hat{h}_{1},\dots,\hat{h}_{\!M}$ are the eigenfunctions associated to the eigenvalues $\hat{\zeta}_{1}\geq\cdots\geq\hat{\zeta}_{M}\geq 0$ of $\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}.$ Further, $\hat{B}_{[i]},$ $B_{[i]},$ $\hat{\Psi}^{\prime\prime\prime}_{[i]},$ $\Psi^{\prime\prime\prime}_{[i]}\in\mathcal{S}_{\mathcal{H}^{\max(i,q)},\mathcal{H}}$ are for each $i,q$ defined by
+where ^(^^$\boldsymbol{\hat{\prod}}^{\dagger}\coloneqq(\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}+\hskip-0.50003pt\gamma_{N}\hskip 0.50003pt\mathbb{I})^{\hskip-0.40974pt-1},$ $(M_{N})\subset\mathbb{N}$ and $(\gamma_{N})\subset(0,\infty)$ are sequences with $M_{N}\to\infty$ and $\gamma_{N}\rightarrow 0,$ and $\hat{h}_{1},\dots,\hat{h}_{\!M}$ are the eigenfunctions associated to the eigenvalues $\hat{\zeta}_{1}\geq\cdots\geq\hat{\zeta}_{M}\geq 0$ of $\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}.$ Further, $\hat{B}_{[i]},$ $B_{[i]},$ $\hat{\Psi}^{\prime\prime\prime}_{[i]},$ $\Psi^{\prime\prime\prime}_{[i]}\in\mathcal{S}_{\mathcal{H}^{\max(i,q)},\mathcal{H}}$ are for each $i,q$ defined by
 $\hat{B}_{[i]}\coloneqq\begin{cases}\,\hat{B}_{i},&\textnormal{if}\;\,1\leq i<q,\\ \,\hat{B}_{q},&\textnormal{if}\;\,i\geq q,\\ \end{cases}\qquad\text{and}\qquad B_{[i]}\coloneqq\begin{cases}\,B_{i},&\textnormal{if}\;\,1\leq i<q,\\ \,B_{q},&\textnormal{if}\;\,i\geq q,\\ \end{cases}$ (5.12)
@@
 The operators $\alpha_{1},\dots,\alpha_{p},\beta_{1},\dots,\beta_{q}\in\mathcal{S_{H}}$ are H-S, and for some $\gamma>0,$ it holds
-$\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle B_{q},h_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}\,(1+i^{2\gamma}\!+j^{2\gamma})<\infty,$
+$\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle B_{q},h_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}\,(1+i^{2\gamma}\!+\hskip-0.50003ptj^{2\gamma})<\infty,$
 where $h_{i}$ and $d_{j}$ denote the eigenfunctions of $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}\colon\mathcal{H}^{q}\rightarrow\mathcal{H}^{q}$ and $\mathscr{C}_{\!\boldsymbol{X}}\colon\mathcal{H}\rightarrow\mathcal{H},$ respectively.
 Assumption 5.6.
-The dimensions of all eigenspaces of $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}$ are bounded above by some $d\in\mathbb{N}.$
-Assumption 5.8.
-For all $N$ sufficiently large, with probability one the eigenvalues of $\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}$ satisfy $\hat{\zeta}_{j}\neq\hat{\zeta}_{j+1}$ for all $j\in\{1,...,M\}.$
+There exists $d\in\mathbb{N}$ such that:
+  • (a)
+    The dimensions of all eigenspaces of $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}$ are bounded above by $d;$
+  • (b)
+    For all $N$ sufficiently large, with probability one the eigenvalues of $\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}$ satisfy $\hat{\zeta}_{j}\neq\hat{\zeta}_{j+1}$ for all $j\in\{1,...,M+d\}.$
 Theorem 5.2.
-Suppose $(X_{k})$ is the fARMA$(p,q)$ process in (5.1).
-Further, let the conditions of Theorem 3.1 and Assumptions 5.55.8 hold, and let $M=M_{\!N}\to\infty$ with $M^{\gamma+1/2}Z^{2}_{\!M}=\mathrm{O}(K^{\beta})$ and $\gamma_{N}=\mathrm{O}(\zeta_{M}M^{-\gamma})$, where $Z_{\!M}\coloneqq(\zeta_{M}-\zeta_{M_{b}})^{-1}$ as in (3.11), with eigenvalues $\zeta_{j}$ of $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}$.
-Then,
+[…]
 $\max_{1\leq i\leq p,1\leq j\leq q}\!\big{\{}\|\hat{\alpha}_{i}-\alpha_{i}\|_{\mathcal{S}},\|\hat{\beta}_{j}-\beta_{j}\|_{\mathcal{S}}\big{\}}$ $=\mathrm{O}_{\operatorname{\mathds{P}}}(M^{-\gamma}).$
@@
 As noted by Granger and Andersen (1978), certain functional time series, such as bilinear processes in separable Hilbert spaces (Hörmann and Kokoszka, 2010), satisfy mild but not standard invertibility.
-Others, like the non-linear MA$(1)$ process $X_{k}=\alpha_{1}(\varepsilon^{2}_{k-1})+\varepsilon_{k}$ in $L^{2}[0,1]$, may be stationary yet fail to be invertible in both senses.
+Others, like the non-linear MA(1)(1)( 1 ) process $X_{k}=\alpha_{1}(\varepsilon^{2}_{k-1})+\varepsilon_{k}$ in $L^{2}[0,1]$, may be stationary yet fail to be invertible in both senses.
 A natural direction for future research is extending estimation methods to mild Granger-Andersen invertibility under suitable assumptions.
 Further, developing a general estimation framework for Banach spaces, potentially even metric spaces, remains an open and promising challenge.
-Acknowledgements
-The authors are grateful to the two referees and the associate editor for their constructive comments on an earlier version of this article, which greatly improved the quality of the paper.
-Parts of the paper were written while Sebastian Kühnert was employed at the University of California, Davis.
-Funding
-Sebastian Kühnert was partially supported by TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (Project number 520388526) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).
-Gregory Rice was partially supported by the Discovery Grant RGPIN 50503-11525 3100 105 from the Natural Science and Engineering Research Council of Canada.
-Appendix A Proofs
+Appendix A Appendix
 Various proofs utilize the operator-valued Hölder inequality for Schatten class operators, along with several other standard results in operator theory (Weidmann, 1980, section 7.1).
@@
 $\|AB\|_{r}\leq\|A\|_{p}\|B\|_{q}.$
+Proof.
+See Weidmann (1980), Theorem 7.8. ∎
+The following lemma follows from elementary norm inequalities and the details are omitted.
 Lemma A.2.
@@
     $\|\boldsymbol{S}\|^{2}_{\mathcal{S}}=\sum^{m}_{i=1}\sum^{n}_{j=1}\|S_{ij}\|^{2}_{\mathcal{S}}.$
-Proof.
-See Weidmann (1980), Theorem 7.8. ∎
 In several places, we make use of the following upper bounds.
@@
 $\|\mathscr{C}^{h}_{\!\boldsymbol{X}^{[L]},\boldsymbol{X}}\|_{\mathcal{N}}\leq\sqrt{L}\operatorname{\mathds{E}}\!\|X_{0}\|^{2},\quad h\in\mathbb{Z}.$ (A.2)
-A.1 Proofs of results in Section 3
 Proof of Proposition 3.1.
@@
 According to the definition of the operator norm, for all $K,L,N$, it holds
-$\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}=\Big{\|}(\hat{\mathscr{C}}+\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{-1}\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}=\sup_{j\leq K}(\hat{\lambda}_{j}+\theta_{\!N})^{-1}=(\hat{\lambda}_{K}+\theta_{\!N})^{-1},$ (A.6)
-$\Big{\|}\mathscr{C}^{\dagger}\!\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}=\Big{\|}(\mathscr{C}+\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{-1}\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}=\sup_{j\leq K}(\lambda_{j}+\theta_{\!N})^{-1}=(\lambda_{K}+\theta_{\!N})^{-1}.$ (A.7)
+$\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}=\Big{\|}(\hat{\mathscr{C}}+\hskip-0.50003pt\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{\hskip-0.40974pt-1}\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}=\sup_{j\leq K}(\hat{\lambda}_{j}+\theta_{\!N})^{-1}=(\hat{\lambda}_{K}+\theta_{\!N})^{-1},$ (A.6)
+$\Big{\|}\mathscr{C}^{\dagger}\!\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}=\Big{\|}(\mathscr{C}+\hskip-0.50003pt\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{\hskip-0.40974pt-1}\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}=\sup_{j\leq K}(\lambda_{j}+\theta_{\!N})^{-1}=(\lambda_{K}+\theta_{\!N})^{-1}.$ (A.7)
 With $K_{b}$ in the definition of $\Lambda_{K}=(\lambda_{K}-\lambda_{K_{b}})^{-1}$ in (3.11), we have
@@
 where the series on the right is finite by Assumption 3.2.
-Finally, from the definition of the H-S norm, with $\Psi_{\!L}=\Psi_{\!L}(K)$ in Assumption A.1, $\mathscr{C}^{\ddagger}=\mathscr{C}(\mathscr{C}+\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{-1},$ and the Kronecker delta $\delta_{jk},$
+Finally, from the definition of the H-S norm, with $\Psi_{\!L}=\Psi_{\!L}(K)$ in Assumption A.1, $\mathscr{C}^{\ddagger}=\mathscr{C}(\mathscr{C}+\hskip-0.50003pt\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{\hskip-0.40974pt-1},$ and the Kronecker delta $\delta_{jk},$
 […]
@@
 $\lim_{L\rightarrow\infty}\!\|\Psi_{\!L}(K)\|^{2}_{\mathcal{S}}\leq K^{2}\sup_{i,j,L}p^{2}_{ijL}<\infty.$ (A.13)
-Altogether, plugging (A.6)–(A.7) and (A.1)–(A.12) in (A.5), and using $\|\mathscr{D}\|_{\mathcal{S}}\leq\sqrt{L}\operatorname{\mathds{E}}\!\|X_{0}\|^{2}$ by (A.2), $(\lambda_{K}+\theta_{\!N})^{-1}\leq\lambda^{-1}_{K}$ and $(\hat{\lambda}_{K}+\theta_{\!N})^{-1}\leq\hat{\Lambda}_{K_{b}-1},$ for each $K,L,N,$ we obtain
+Altogether, plugging (A.6)–(A.7) and (A)–(A.12) in (A.5), and using $\|\mathscr{D}\|_{\mathcal{S}}\leq\sqrt{L}\operatorname{\mathds{E}}\!\|X_{0}\|^{2}$ by (A.2), $(\lambda_{K}+\theta_{\!N})^{-1}\leq\lambda^{-1}_{K}$ and $(\hat{\lambda}_{K}+\theta_{\!N})^{-1}\leq\hat{\Lambda}_{K_{b}-1},$ for each $K,L,N,$ we obtain
 $\|\hat{\Psi}_{\!L}-\Psi_{\!L}(K)\|_{\mathcal{S}}$ $\leq\hat{\Lambda}_{K_{b}-1}\|\hat{\mathscr{D}}-\mathscr{D}\|_{\mathcal{S}}+2\operatorname{\mathds{E}}\!\|X_{0}\|^{2}\hat{\Lambda}_{K_{b}-1}\sqrt{(K+dL)L}\,\|\hat{\mathscr{C}}-\mathscr{C}\|_{\mathcal{L}}\,(\hat{\Lambda}_{K_{b}-1}+2\Lambda_{K}+1)$
@@
 ∎
-A.2 Proofs of results in Section 4
 Proof of Proposition 4.1.
@@
 ∎
-A.3 Proofs of results in Section 5
 Proof of Corollary 5.1.
@@
 Follows from Proposition 4.2 by putting $\beta_{j}=\phi_{j}$ and $\hat{\beta}_{j}=\hat{\phi}_{j}.$
-To derive the asymptotic result for our fARMA(1,1)11(1,1)( 1 , 1 ) operators, we derive a preliminary result based on the following assumption.
+To derive the asymptotic result for our fARMA(1,1)(1,1)( 1 , 1 ) operators, we derive a preliminary result based on the following assumption.
 Assumption A.2.
@@
 Lemma A.4.
-Suppose $(X_{k})$ is the fARMA(1,1)11(1,1)( 1 , 1 ) process in (5.2). Further, let the conditions of Theorem 3.1, Assumptions 5.1, 5.35.4, A.2, and $\gamma_{N}=\mathrm{O}(K^{-\beta})$ hold. Then,
+Suppose $(X_{k})$ is the fARMA(1,1)(1,1)( 1 , 1 ) process in (5.2). Further, let the conditions of Theorem 3.1, Assumptions 5.1, 5.3, A.2, and $\gamma_{N}=\mathrm{O}(K^{-\beta})$ hold. Then,
 $\max\!\big{\{}\|\hat{\alpha}_{1}-\alpha_{1}(M)\|_{\mathcal{S}},\|\hat{\beta}_{1}-\beta_{1}(M)\|_{\mathcal{S}}\big{\}}$ $=\mathrm{O}_{\operatorname{\mathds{P}}}\big{(}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}\big{)}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta}).$
@@
 Lemma A.5.
-Let $(X_{k})$ be the fARMA$(p,q)$ process in (5.1). Moreover, let the conditions of Theorem 3.1, Assumptions 5.5, 5.75.8, A.3 and $\gamma_{N}=\mathrm{O}(K^{-\beta})$ hold. Then,
+[…]
 $\max_{1\leq i\leq p,1\leq j\leq q}\!\big{\{}\|\hat{\alpha}_{i}-\alpha_{i}(M)\|_{\mathcal{S}},\|\hat{\beta}_{j}-\beta_{j}(M)\|_{\mathcal{S}}\big{\}}=\mathrm{O}_{\operatorname{\mathds{P}}}\big{(}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}\big{)}=\mathrm{O}_{\operatorname{\mathds{P}}}(K^{-\beta}).$
@@
 $\hat{B}_{q}-B_{q}(M)$ $=\big{(}\Psi^{\prime}_{[q]}-\hat{\Psi}^{\prime}_{[q]}\big{)}\boldsymbol{\hat{\prod}}^{\dagger}\coprod^{\hat{h}_{\!M}}_{\hat{h}_{1}}+\,\Psi^{\prime}_{[q]}\Big{(}\boldsymbol{\prod}^{\dagger}\!\coprod^{h_{\!M}}_{h_{1}}-\boldsymbol{\hat{\prod}}^{\dagger}\!\coprod^{\hat{h}_{\!M}}_{\hat{h}_{1}}\Big{)}+B_{q}\Big{(}\boldsymbol{\prod}^{\ddagger}\!\coprod^{h_{\!M}}_{h_{1}}-\,\mathbb{I}\Big{)}.$
-Further, as in the proof of Lemma A.4, together with (A.20) and Assumptions 5.75.8, it holds
+Further, as in the proof of Lemma A.4, together with (A.20) and Assumption 5.6, it holds
 $\|\hat{B}_{q}-B_{q}(M)\|_{\mathcal{S}}$ $\leq(\hat{\zeta}_{M}+\gamma_{N})^{-1/2}\|\hat{\Psi}_{\!L}-\Psi_{\!L}\|_{\mathcal{S}}+\|\Psi^{\prime}_{[q]}\|_{\mathcal{S}}\Big{\|}\boldsymbol{\prod}^{\dagger}\!\coprod^{h_{\!M}}_{h_{1}}-\boldsymbol{\hat{\prod}}^{\dagger}\!\coprod^{\hat{h}_{\!M}}_{\hat{h}_{1}}\Big{\|}_{\mathcal{L}}+\gamma_{N}\zeta^{-1}_{M}\|B_{q}(M)\|_{\mathcal{S}}$
@@
 Follows from the argumentation in the proofs of Theorem 5.1 and Lemma A.5 and the requirements for our sequences. ∎
+Acknowledgements
+The authors are grateful to the editor, the associate editor and the two referees for their constructive comments on an earlier version of this article, which greatly improved the quality of the paper.
+Parts of the paper were written while Sebastian Kühnert was employed at the University of California, Davis.
+Funding
+Alexander Aue was partially supported by NSF DMS 2515821.
+Sebastian Kühnert was partially supported by TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (Project number 520388526) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).
+Gregory Rice was partially supported by the Discovery Grant RGPIN 50503-11525 3100 105 from the Natural Science and Engineering Research Council of Canada.